### v4

# Seth Pettie

Abstract
In this work we uncover an intimate relation between Lévy processes and data sketches for generalized moment estimation and weighted sampling. Let $\mathbf{x}\in(\mathbb{R}^{d})^{n}$ be a vector subject to element updates and $f:\mathbb{R}^{d}\to\mathbb{R}_{+}$. By definition, the $f$-moment of $\mathbf{x}$ is
$f(\mathbf{x})=\sum_{v\in[n]}f(\mathbf{x}(v)).$
The $f$-moment estimation problem is to $(1\pm\epsilon)$-approximate $f(\mathbf{x})$ whereas the $f$-sampling problem is to select an index $v_{*}\in[n]$ with probability $f(\mathbf{x}(v_{*}))/f(\mathbf{x})$. Our primary conceptual contributions are
  • I.
    to draw a close connection between $f$-moment estimation sketches of $\mathbf{x}\in(\mathbb{R}^{d})^{n}$ in the turnstile model and generic Lévy processes over $\mathbb{R}^{d}$, and
  • II.
    to draw a close connection between $f$-samplers of $\mathbf{x}\in\mathbb{R}^{n}$ in the incremental model (positive updates only) and one-dimensional, non-negative Lévy processes, aka subordinators.
Through these connections, we can apply the powerful Lévy-Khintchine representation theorem from the theory of Lévy processes to design new sketches for $f$-moment estimation and sampling. Out technical results are as follows.
  • We give a systematic method for transforming any Lévy process $X=(X_{t})_{t\geq 0}$ in $\mathbb{R}^{d}$ into an $O(\epsilon^{-2}\log^{2}n)$-bit sketch that estimates the $f_{X}$-moment, where $f_{X}(z)=-\log\mathbb{E}e^{i\langle z,X_{1}\rangle}$ is the characteristic exponent of $X$. This method handles essentially all known $f$-moments that can be estimated with $\operatorname{poly}(\epsilon^{-1},\log n)$-size sketches [1, 29, 18, 37, 28, 8, 54] in a uniform way, broadens the class of tractable functions, and allows us to estimate multivariate functions, when $d>1$.
  • In the one-dimensional incremental setting, we transform any non-negative Lévy process (a subordinator) $X$ into a $G_{X}$-sampler, where $G_{X}(z)=-\log\mathbb{E}e^{-zX_{1}}$ is the Laplace exponent of $X$. These samplers require essentially minimal space, sample with precisely correct probabilities, and have zero probability of error. They are distinguished from recent work on $G$-samplers [16, 33], which either introduce $(1\pm\epsilon)$-approximation in the probabilities, a non-zero failure probability, or additional space.
1 Introduction
The study of sketches and the streaming model dates back to the late 1970s, whose early work included Morris’s [43] approximate counter, Munro and Patterson’s [44] selection algorithms, and the Boyer-Moore MJRTY algorithm [6]. In 1983 Flajolet and Martin [24] designed and analyzed the first “modern” data sketch called PCSA, [^1] Probabilistic Counting with Stochastic Averaging which estimates the cardinality of a set. PCSA works only in incremental streams whereas all linear sketches, such as the celebrated AMS sketch of Alon, Matias, and Szegedy [1] for estimating $F_{2}$ moments, operate on a vector of elements subject to both increments and decrements. After decades of intense investigation into streaming and sketching we now have a good—but still incomplete—understanding of which statistics are tractable, meaning they are approximable with polylogarithmic-size sketches [9, 7, 8]. The purpose of this paper is to approach this tractability question from a new direction and find answers that are not merely true but have significant explanatory power.
We shall begin by defining a single algebraic streaming model that captures the incremental model, turnstile model, and others.
Definition 1 ($M$-turnstile model [53]).
Let $(M,+)$ be a commutative monoid with identity 0. The $M$-turnstile model is defined as follows. Let $n]=\{1,2,\ldots,n\}$ be the universe. The state vector $\mathbf{x}=(\mathbf{x}(1),\ldots,\mathbf{x}(n))\in M^{n}$ is initialized as $0^{n}$ and gets updated by a stream of pairs in the form of $(v,y)$, where $v\in[n]$ and $y\in M$.
  • $\mathsf{Update}(v,y)$ : $\mathbf{x}(v)\leftarrow\mathbf{x}(v)+y$.
The assumption that $M$ is commutative and associative allows the updates to be collected in a distributed system. The point of the $M$-turnstile model is to identify algorithmic functionality with the mathematical structure of the monoid $(M,+)$. For example, if $M$ is idempotent, then duplicated updates are ignored; if $M$ is a group, then “deletions” are allowed; if $M$ is a continuous space, then fractional updates are allowed; if $M$ is multi-dimensional, then attribute-wise updates are allowed. The usual integer turnstile model corresponds to $\mathbb{Z}$-turnstile while the incremental setting corresponds to $\mathbb{N}$-turnstile. In this paper we consider 1-dimensional turnstiles ($\mathbb{R}$ or $\mathbb{Z}$), multidimensional turnstiles ($\mathbb{R}^{d}$), and incremental turnstiles ($\mathbb{R}_{+}$ or $\mathbb{N}$).
Let us recall three generic streaming problems. Here $\mathbb{R}_{+}$ is the set of non-negative reals.
Problem 1 ($f$-moment estimation in the $\mathbb{R}^{d}$-turnstile model).
Fix $d\in\mathbb{Z}_{+}$. Let $f:\mathbb{R}^{d}\to\mathbb{C}$ be a function with $f(0)=0$, and let $\mathbf{x}\in(\mathbb{R}^{d})^{n}$ be a vector subject to streaming updates to its coordinates. The problem is to estimate the $f$-moment $f(\mathbf{x})$, where
$f(\mathbf{x})\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\sum_{v\in[n]}f(\mathbf{x}(v)).$
Problem˜2 is a special case of Problem˜1, but it is nonetheless useful to highlight as a separate problem, due to its connection to Problem˜3 and a separate suite of techniques.
Problem 2 ($G$-moment estimation in the $\mathbb{R}_{+}$-turnstile model).
Let $G:\mathbb{R}_{+}\to\mathbb{R}_{+}$ with $G(0)=0$. Let $\mathbf{x}\in\mathbb{R}_{+}^{n}$ be the state vector of the current stream. Estimate the $G$-moment $G(\mathbf{x})\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\sum_{v\in[n]}G(\mathbf{x}(v))$ over the stream.
Problem 3 ($G$-sampler in the $\mathbb{R}_{+}$-turnstile model).
Let $G:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a function with $G(0)=0$ and let $\mathbf{x}\in\mathbb{R}_{+}^{n}$ be the vector subject to streaming updates. Return a $G$-sample $u\in[n]$ with probability $G(\mathbf{x}(u))/G(\mathbf{x})$.
These three problems encompass a large body of research in the streaming/sketching literature. We give a brief overview of the prior research on these problems.
$f$-moments in the $\mathbb{R}$- and $\mathbb{Z}$-turnstile.
Alon, Matias, and Szegedy [1] estimate the $F_{2}$-moment ($f(x)=x^{2}$) with a data structure now commonly known as the AMS sketch. Indyk [29] designed a class of sketches for estimating the $F_{p}$ moments, with $p\in(0,2]$. When $p>2$, Bar-Yossef, Jayram, Kumar, and Sivakumar [4] proved that estimating the $F_{p}$-moment requires $\Omega(\operatorname{poly}(n))$ space. The $F_{0}$-moment ($f(x)=\mathbbm{1}\!\left\{{x\neq 0}\right\}$) has been estimated in three distinct ways. [^2] Here $\mathbbm{1}\!\left\{{P}\right\}$ is the indicator for $P$, i.e., 1 if the predicate $P$ is true, and 0 otherwise. Cormode, Datar, Indyk, and Muthukrishnan [18] approximate $F_{0}$ by $F_{\alpha}$ with very small $\alpha>0$. Kane, Nelson, and Woodruff [37] project each element onto $\mathbb{Z}_{p}$, $p>\epsilon^{-1}\log n$ being a random prime, which effectively reduces $F_{0}$-estimation to cardinality estimation. Very recently, Wang [54] estimates $F_{0}$ by estimating the underlying harmonic components separately and then summing them up. In the $\mathbb{R}^{d}$-turnstile model, the $F_{p,q}$ hybrid moment is defined by $f(x_{1},\ldots,x_{d})=(\sum_{j=1}^{d}|x_{j}|^{p})^{q}$. Ganguly, Bansal, and Dube [28] approximate $F_{p,q}$ for $p\in(0,2],q\in(0,1]$ in $\tilde{O}(\epsilon^{-2})$ space, and Jayram and Woodruff [34] give polynomial space bounds on the complexity of $F_{p,q}$ estimation for arbitrary values of $p,q$.
Braverman and Ostrovsky [9] considered the problem of characterizing the class of tractable functions $f$ for which $f$-moments could be approximated to within a $1\pm\epsilon$ factor in $\operatorname{poly}(\epsilon^{-1},\log n)$ space. They managed to characterize all functions $f\colon\mathbb{Z}\to\mathbb{R}_{+}$ that are symmetric (f(x)=f(x)f(x)=f(-x)) and increasing on $0,\infty)$. Braverman, Chestnut, Woodruff, and Yang [8] extended the characterization to almost all symmetric functions, excluding a class they termed “nearly periodic functions.”
$G$-moments in $\mathbb{R}_{+}$-turnstile.
The incremental setting ($\mathbb{N}$-turnstile or $\mathbb{R}_{+}$-turnstile) leads to different sketching techniques and different space lower bounds. Flajolet and Martin’s [24] PCSA sketch estimates the cardinality (number of distinct elements) $\|\mathbf{x}\|_{0}$, which corresponds to the $G$-moment with $G(x)=\mathbbm{1}\!\left\{{x\neq 0}\right\}$. Flajolet, Fusy, Gandouet, and Meunier’s [23] HyperLogLog sketch is the most widely deployed sketch for cardinality estimation. The most efficient sketches in terms of space vs. estimation error [47, 55, 38, 3] are based on entropy-compressed versions of PCSA with optimum estimators. The cardinality sketches above are analyzed in the random oracle model, where one can evaluate uniformly random hash functions. When the sketch stores its own hash functions, Kane, Nelson, and Woodruff [37] and Błasiok [5] designed cardinality sketches with an $(\epsilon,\delta)$-guarantee meeting the $\Omega(\log n+\epsilon^{-2}\log\delta^{-1})$ lower bound [30, 35, 1], up to a large constant factor.
The problem of $F_{p}$-moment estimation ($G(x)=x^{p}$) can be solved by Indyk’s sketches for $p\in(0,2]$. Nevertheless, there are more efficient $F_{p}$-moment sketches in the incremental setting when $p\in(0,1)$; see Cohen [14] and Jayaram and Woodruff [32]. Cohen [14] estimates the $G$-moment, where $G$ is in the class of “soft concave sublinear functions,” which are intended to approximate cap-statistics of the form $G(x)=\min\{x,C\}$.
$G$-sampling in $\mathbb{R}_{+}$-turnstile.
Vitter’s [52] reservoir sampling can be regarded as a $G$-sampler in the $\mathbb{N}$-turnstile model, where $G(x)=x$, i.e., elements are sampled proportional to their counts. Cohen’s [13] Min-Sampler hashes elements and samples the index $v$, $\mathbf{x}(v)>0$, having the smallest hash value. The Min-Sampler is insensitive to duplicates and thus solves the $G$-sampling problem with $G(x)=\mathbbm{1}\!\left\{{x>0}\right\}$ (disinct sampling).
For generic $G$-sampling, Cohen and Geri [16] convert the $G$-estimators in [14] into approximate $G$-samplers for soft concave sublinear functions. Generic $G$-samplers with precisely correct sampling probabilities have only been studied recently, by Jayaram, Woodruff, and Zhou [33], where they combine reservoir sampling and rejection sampling to ensure correct sampling probabilities, conditioned on a sample being returned.
1.1 A New Perspective
The main take-away message of the present work is that
$f$-moment estimation in the $\mathbb{R}^{d}$-turnstile model and $G$-moment estimation and $G$-sampling in the $\mathbb{R}_{+}$-turnstile model can all be done in a uniform manner by simulating Lévy processes.
Lévy processes have been studied since the early 20th century, and have been used to model phenomena in various fields, e.g., physics (how does a gas particle move?) and finance (how does the stock price change?). We will show that all Lévy processes have algorithmic interpretations in the context of streaming sketches, and the fundamental Lévy-Khintchine representation theorem leads to a unified view of sketching for moment estimation and sampling. To our knowledge, this is the first application of the Lévy-Khintchine theorem in theoretical computer science.
We give a detailed technical synopsis of Lévy processes in Section˜2. For the time being, a Lévy process $(X_{t})_{t\geq 0}$, where $X_{t}\in\mathbb{R}^{d}$, is defined by having independent, stationary increments. I.e., for any $t_{1},t_{2}\in\mathbb{R}_{+}$, $(X_{t_{1}+t_{2}}-X_{t_{1}})\sim X_{t_{2}}$, and for any $t_{1}<t_{2}<\cdots<t_{k}$, the increments $X_{t_{1}},X_{t_{2}}-X_{t_{1}},\ldots,X_{t_{k}}-X_{t_{k-1}}$ are mutually independent. Some common one-dimensional Lévy processes are
Linear drift.
$X_{t}=\gamma t$ with drift rate $\gamma$.
Wiener process/Brownian motion.
$X_{t}\sim\mathcal{N}(0,t\sigma^{2})$ with variance $\sigma^{2}$.
Poisson process.
$X_{t}\sim\mathrm{Poisson}(\lambda t)$ with rate $\lambda$.
Compound Poisson process.
$X_{t}=\sum_{k=1}^{Z}J_{k}$, where the number of jumps $Z\sim\mathrm{Poisson}(\lambda t)$ follows a Poisson process, while each jump $J_{k}\sim\mathcal{J}$ is sampled independently from a common jump distribution.
$\alpha$-stable process.
Defined by $X_{t}/t^{1/\alpha}\sim X_{1}\sim\text{$\alpha$-stable}$. (For $\alpha\in(0,2]$, the $\alpha$-stable random variable $X_{1}$ on $\mathbb{R}$ is defined by $\mathbb{E}e^{izX_{1}}=e^{-|z|^{\alpha}}$, for any $z\in\mathbb{R}$.)
The Lévy processes defined above can all be generalized to higher dimensions. They may also be combined, as Lévy processes are closed under addition.
The characteristic function $\varphi_{X}(z)=\mathbb{E}e^{izX}$ of a random variable $X\in\mathbb{R}$ is equivalent to the Fourier transform of its pdf, when its pdf exists. It uniquely determines the distribution of $X$. For example, whenever it exists, the $k$th moment of $X$ can be recovered from the $k$th derivative of $\varphi_{X}$ evaluated at zero. When $X\in\mathbb{R}^{d}$, $\varphi_{X}(z)=\mathbb{E}e^{i\langle X,z\rangle}$, where $z\in\mathbb{R}^{d}$ and $\langle X,z\rangle$ is the inner product.
The Lévy-Khintchine representation theorem identifies every Lévy process $X$ with its characteristic exponent $f_{X}\colon\mathbb{R}^{d}\to\mathbb{C}$, where for any $z\in\mathbb{R}^{d}$,
$\mathbb{E}\exp(i\langle X_{t},z\rangle)=\exp(-tf_{X}(z)).$
When $X$ is a non-negative Lévy process in one dimension (also known as a subordinator), it is characterized by its Laplace exponent $G_{X}\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$, where for any $z\in\mathbb{R}_{+}$,
$\mathbb{E}\exp(-zX_{t})=\exp(-tG_{X}(z)).$
Note that the stationary increments property implies that a non-negative Lévy process has only non-negative increments.
Roadmap.
In the remainder of the introduction we give a lightly technical introduction to how Lévy processes naturally arise in the study of generalized moment estimation and sampling problems. In Section˜1.2 we show how linear sketches [1, 29] explicitly or implicitly use infinitely divisible distributions and Lévy processes. In Section˜1.3 we show a natural connection between the $f$-moment estimation problem, Lévy processes, and the Lévy-Khintchine representation theorem, paving the way to answer Problem˜1 in turnstile streams. Section˜1.4 illustrates a connection between non-negative Lévy processes (subordinators) and incremental streams, which allows one to address both moment estimation (Problem˜2) and sampling (Problem˜3) in incremental streams. In Section˜1.5 we give a formal statement of our new results, and in Section˜1.6 we lay out the organization of the remainder of the paper.
1.2 Lévy Processes and Linear Sketching
We briefly explain why Lévy processes naturally lie at the heart of linear sketching. Suppose $X$ is the random memory state of a linear sketch [^3] It suffices to consider linear sketches in the turnstile model [40]. of a particular input vector $\mathbf{x}\in\mathbb{R}^{n}$, with $\hat{f}(X)$ being an estimate of the $f$-moment $f(\mathbf{x})$. Now consider the situation where the input is replicated $w$ times over disjoint domains, i.e., $\mathbf{x}_{w}\in\mathbb{R}^{nw}$, with $\mathbf{x}_{w}(v)=\mathbf{x}(v\operatorname{mod}n)$. By construction, $f(\mathbf{x}_{w})=w\cdot f(\mathbf{x})$ and the estimate becomes $\hat{f}(X^{(1)}+\cdots+X^{(w)})$ where $X^{(j)}$ are i.i.d. copies of $X$. In other words, we should have
$\frac{\hat{f}(X^{(1)}+\cdots+X^{(w)})}{w}\approx f(\mathbf{x}),$
for any $w\in\mathbb{Z}_{+}$ and also for the limiting case as $w\to\infty$. Therefore, no matter how complicated the distribution of the linear sketch may be, the sum $X^{(1)}+\cdots+X^{(w)}$ will converge to some well-behaved limiting distribution, with a proper normalization depending on $X$ and $w$. For the AMS sketch [1], $X$ happens to be sub-Gaussian and therefore the normalized sum converges to a Gaussian. For Indyk’s [29] stable sketch, $X$ is $\alpha$-stable, and therefore the normalized sum remains $\alpha$-stable. These are merely two special cases of Lévy processes. In general, if $X=(X_{t})_{t\geq 0}$ is a Lévy process, then by the stationary and independent increments properties, $X^{(1)}+\cdots+X^{(w)}=(X^{(1)}_{t}+\cdots+X^{(w)}_{t})_{t\geq 0}\sim(X_{wt})_{t\geq 0}$. In other words, summing i.i.d. Lévy processes is equivalent to simply rescaling time. After normalizing the time scale, [^4] To see how this “time normalization” generalizes the typical scalar normalization of stable variables, note that for $\alpha$-stable variables, the normalization is $w^{-1/\alpha}(X^{(1)}+\cdots+X^{(w)})$, which is equivalent to scale the time down by $w$ since $w^{-1/\alpha}X_{wt}\sim X_{t}$ if $(X_{t})_{t\geq 0}$ is $\alpha$-stable. we see Lévy processes are stable under i.i.d. sums. Lévy processes therefore form a mathematical closure of linear sketches in terms of their limiting distributions.
In practice it usually suffices to construct some algorithmically simple random projection in the domain of attraction of the limiting process. For example, the AMS sketch for estimating $F_{2}$ does not need to explicitly use Gaussians in its projection; Rademacher ($\{-1,1\}$) random variables suffice. For a prototypical solution, it is convenient to consider linear sketches with projection sampled directly from Lévy processes. The distribution of the sketch will always lie in the space of Lévy processes and is therefore easier to track.
We now demonstrate how to convert Lévy processes to streaming sketches. Throughout the paper we work in the random oracle model, in which we can evaluate uniformly random hash functions $H\colon\mathbb{Z}\to[0,1]$. This is not a limiting assumption as it can be removed in a black-box way by using pseudorandom generators against space bounded computation, at a small loss in space efficiency; see [29, 45].
1.3 $f$-Moment Estimation and Lévy Processes
Given any Lévy process $X=(X_{t})_{t\geq 0}$ on $\mathbb{R}^{d}$, by the Lévy-Khintchine representation theorem (Theorem˜6 in Section˜2), there exists a function $f=f_{X}:\mathbb{R}^{d}\to\mathbb{C}$ such that for any $t\geq 0$ and $z\in\mathbb{R}^{d}$, $\mathbb{E}e^{i\langle z,X_{t}\rangle}=e^{-tf(z)}$. Suppose now we project the input vector to a single cell $C_{t}$ by sampling the Lévy processes at time $t$, as follows.
$C_{t}$ $=\sum_{v\in[n]}\left\langle X_{t}^{(v)},\mathbf{x}(v)\right\rangle.$
Here the $X_{t}^{(v)}$ are i.i.d. copies of the Lévy process $X_{t}$. Clearly $C_{t}$ is a linear sketch that can be maintained over a distributed stream. We thus have
$\mathbb{E}e^{iC_{t}}$ $=\mathbb{E}e^{i\sum_{v\in[n]}\left\langle X_{t}^{(v)},\mathbf{x}(v)\right\rangle}$ (definition of $C_{t}$)
$=\prod_{v\in[n]}\mathbb{E}e^{i\left\langle X_{t}^{(v)},\mathbf{x}(v)\right\rangle}$ (by independence)
$=\prod_{v\in[n]}e^{-tf(\mathbf{x}(v))}$ (by Lévy-Khintchine)
$=e^{-tf(\mathbf{x})},$ (definition of $f(\mathbf{x})=\sum_{v\in[n]}f(\mathbf{x}(v))$)
from which the $f$-moment $f(\mathbf{x})$ can be recovered by choosing a suitable time $t\approx\Theta(1/|f(\mathbf{x})|)$. Of course, we do not know $f(\mathbf{x})$ in advance and therefore do not know the optimum $t$ in advance, but we achieve reasonable coverage by maintaining $(C_{t})$ for many $t$, evenly spaced on a logarithmic scale. This is the key observation that lets us estimate any $f$-moment, so long as $f$ is the characteristic exponent of some Lévy process.
1.4 $G$-Sampler, $G$-Moment Estimation, and Subordinators
Perhaps surprisingly, the connection between Lévy processes and streaming sketches goes beyond the linear case. We now demonstrate that the min-based sketches/samplers are closely related to non-negative Lévy processes, also known as subordinators. To index the fresh randomness associated with an update, we suppose updates occur at times $k=1,2,\ldots$. The min-based samplers are based on the following generic min sketch.
Definition 2 (generic min sketch).
A generic min sketch is a pair $(v_{*},h_{*})\in([n]\cup\{\perp\})\times(\mathbb{R}_{+}\cup\{\infty\})$ initialized as $(\perp,\infty)$. It is associated with a certain hash function $H:[n]\times\mathbb{R}_{+}\times\mathbb{Z}_{+}\to\mathbb{R}_{+}$ provided by the random oracle. A vector update $\mathbf{x}(v)\leftarrow\mathbf{x}(v)+\Delta$ is handled as follows.
$\mathsf{Update}(v,\Delta)$ (issued at time $k$):
If $H(v,\Delta,k)<h_{*}$, then set $(v_{*},h_{*})\leftarrow(v,H(v,\Delta,k))$.
Assume for simplicity that $\Delta=1$ in all updates. Suppose $\mathbf{x}(v)=w$, being incremented at times $k_{1},\ldots,k_{w}$. The minimum hash value produced at index $v$ is thus
$Z_{v}=\min(H(v,1,k_{1}),\ldots,H(v,1,k_{w})).$
The observation is that, no matter what the weight function $G$ is, the probability that $v$ gets sampled should only depend on $w$, rather than the insertion timestamps $(k_{1},\ldots,k_{w})$. Thus, we want the random sequence
$H(v,1,1),H(v,1,2),H(v,1,3),\ldots$ (1)
to be exchangeable. [^5] A finite random vector $(Y_{1},\ldots,Y_{k})$ is exchangeable if its distribution remains unchanged after reordering the coordinates. An infinite random sequence is exchangeable if any finite prefix is exchangeable. See [10, Definition 2.1] for a recent treatment. De Finetti’s theorem [^6] This is a classic result by De Finetti from 1931 on exchangeable random sequences. We use the version described in [10, Theorem 2.2]. then implies that there is a non-negative random process $X=(X_{t})_{t\geq 0}$ such that
$(H(v,1,k))_{k\in\mathbb{Z}_{+}}=(\{\inf\{t\geq 0:X_{t}>E_{k}\}\})_{k\in\mathbb{Z}_{+}},$
where the $E_{k}\sim\mathrm{Exp}(1)$ are i.i.d. standard exponential random variables. In addition, in order to sample with the correct probability, we need the minimum hash value $\min_{j\in[w]}(H(v,1,k_{j}))$ to distribute as an exponential random variable (also observed in [31]), which will be true if $X$ is a Lévy process. Note that $X$ needs to be non-negative in De Finetti’s theorem. Now suppose $X$ is a subordinator (non-negative Lévy process) with Laplace exponent $G_{X}:\mathbb{R}_{+}\to\mathbb{R}_{+}$. Then for $z>0$,
$\mathbb{P}(\min(H(v,1,k_{1}),\ldots,H(v,1,k_{w}))>z)$
$=\mathbb{P}(\min(H(v,1,1),\ldots,H(v,1,w))>z)$ ((1) is exchangeable)
$=\mathbb{P}(\min(\inf\{t\geq 0:X_{t}>E_{1}\},\ldots,\inf\{t\geq 0:X_{t}>E_{w}\})>z)$ (De Finetti’s theorem)
$=\mathbb{P}(\inf\{t\geq 0:X_{t}>\mathrm{Exp}(w)\}>z)$
$=\mathbb{P}(X_{z}\leq\mathrm{Exp}(w))$ (non-negativity)
$=\mathbb{E}(\mathbb{P}(X_{z}\leq\mathrm{Exp}(w))\mid X_{z})$ (law of total probability)
$=\mathbb{E}e^{-wX_{z}}=e^{-zG_{X}(w)}.$ (Lévy-Khintchine)
Note that $\mathbb{P}(\mathrm{Exp}(\lambda)>z)=e^{-z\lambda}$. Therefore we know that
$\min(H(v,1,k_{1}),\ldots,H(v,1,k_{w}))\sim\mathrm{Exp}(G_{X}(w))=\mathrm{Exp}(G_{X}(\mathbf{x}(v))).$
I.e., the minimum hash value produced by an element $v$ is exactly an exponential random variable with rate $G(\mathbf{x}(v))$. This is the underlying intuition for the Lévy-Min-Sampler studied in Section˜4. The precise simulation of $\mathrm{Exp}(G_{X}(w))$ can be used to emulate existing cardinality estimators with the cardinality of $\mathbf{x}$ ($\|\mathbf{x}\|_{0}=|\mathrm{supp}(\mathbf{x})|$) replaced by the $G$-moment $G(\mathbf{x})$. This is the underlying intuition for the LévyPCSA and LévyHyperLogLog sketches studied in Section˜5.
1.5 New Results
We prove two main theorems. Theorem˜1 connects the Lévy-Khintchine representation theorem for generic Lévy processes with $f$-moment estimation in the $\mathbb{R}^{d}$-turnstile model, while Theorem˜2 connects subordinators with $G$-sampling in the $\mathbb{R}_{+}$-turnstile. Refer to Table˜1 for some standard notation.
Notation Definition Notes
$|x|$ $(\sum_{j=1}^{d}x_{j}^{2})^{1/2}$ $x\in\mathbb{R}^{d}$, Euclidean norm
$|z|$ $(z\bar{z})^{1/2}$ $z\in\mathbb{C}$, modulus
$\|\mathbf{x}\|_{0}$ $\sum_{v\in[n]}\mathbbm{1}\!\left\{{\mathbf{x}(v)\neq 0}\right\}$ $\mathbf{x}\in(\mathbb{R}^{d})^{n}$
$\|\mathbf{x}\|_{\infty}$ $\max_{v\in[n]}|\mathbf{x}(v)|$ $\mathbf{x}\in(\mathbb{R}^{d})^{n}$
$\mathbb{T}$ complex unit circle identified by $0,2\pi)$
$\mathbb{S}_{d-1}$ $\{x\in\mathbb{R}^{d}:|x|=1\}$ unit sphere in $\mathbb{R}^{d}$
$\mathbb{V}Z$ $\mathbb{E}(Z-\mathbb{E}Z)\overline{(Z-\mathbb{E}Z)}$ variance of complex $Z$
Table 1: Notation
Theorem 1 (Lévy-Tower, Section˜3).
Let $f:\mathbb{R}^{d}\to\mathbb{C}$ be any function of the form
$f(z)=\frac{1}{2}\left\langle z,Az\right\rangle-i\left\langle\gamma,z\right\rangle+\int_{\mathbb{R}^{d}}\left(1+i\left\langle z,s\right\rangle\mathbbm{1}\!\left\{{|s|<1}\right\}-e^{i\left\langle z,s\right\rangle}\right)\,\nu(ds),$
where $A$ is a covariance matrix, $\gamma\in\mathbb{R}^{d}$ is a drift, and $\nu$ is a positive measure over $\mathbb{R}^{d}$ with $\int_{\mathbb{R}^{d}}\min\{|s|^{2},1\}\,\nu(ds)<\infty$. The Lévy-Tower, parameterized by $f$, is a mergeable sketch that occupies $O(\epsilon^{-2}\log n)$ words. For any input stream $\mathbf{x}\in(\mathbb{R}^{d})^{n}$ with $|f(\mathbf{x})|\in[1,\operatorname{poly}(n)]$, the Lévy-Tower returns an estimate $\widehat{f(\mathbf{x})}$ that with probability 99/100 satisfies:
$\left|\widehat{f(\mathbf{x})}-f(\mathbf{x})\right|\leq O(\epsilon|f(\mathbf{x})|).$
The Lévy-Tower sketch improves our understanding of the tractability of one-dimension function moments. In particular, it implies the tractability of many nearly periodic functions that were not previously classified. See Section˜7 for a discussion of the new result and the existing tractability results. It also implies the tractability of a large class of multidimensional moments which have not been considered before.
Theorem 2 (Lévy-Min-Sampler, Section˜4.3).
Let $G:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be any function of the form
$G(z)=c\mathbbm{1}\!\left\{{z>0}\right\}+\gamma_{0}z+\int_{0}^{\infty}(1-e^{-zs})\,\nu(ds),$
where $c,\gamma_{0},\geq 0$, and $\nu$ is any positive measure such that $\int_{0}^{\infty}\min\{s,1\}\,\nu(ds)<\infty$. There is a min-based sketch storing only a single pair $(v_{*},h_{*})\in[n]\times\mathbb{R}_{+}$, where $v_{*}\in[n]$ is the sampled index and $h_{*}\in\mathbb{R}_{+}$ is the minimum hash value. For any given input vector $\mathbf{x}\in\mathbb{R}_{+}^{n}$ it is guaranteed that
  • For any $u\in[n]$, $\mathbb{P}(v_{*}=u)=G(\mathbf{x}(u))/G(\mathbf{x})$.
  • $h_{*}\sim\mathrm{Exp}(G(\mathbf{x}))$.
The space required is therefore only 2 words.
The raw version of the Lévy-Min-Sampler (Section˜4.1) is built on the structure theory of min-wise infinitely divisible, exchangeable sequences. See Brück, Mai, and Scherer [10] for a recent treatment. Cohen [16] developed approximate $G$-samplers for soft concave sublinear functions, which are a subset of Laplace exponents. Jayaram, Woodruff, and Zhou [33] have developed $G$-samplers with precisely correct sampling probabilities for functions $G:\mathbb{N}\to\mathbb{R}$ such that $\max_{z\in\mathbb{N}}(G(z)-G(z-1))<\infty$ with $O(\frac{\sum_{v\in[n]}\mathbf{x}(v)}{G(\mathbf{x})}\log n)$ bits of space. For $G$ being a Laplace exponent, the factor $\frac{\sum_{v\in[n]}\mathbf{x}(v)}{G(\mathbf{x})}$ in [33] can be $\Omega(\operatorname{poly}(n))$ when $\sum_{v\in[n]}\mathbf{x}(v)\gg G(\mathbf{x})$. The significance of Theorem˜2 is that the sampling probability is precisely correct and the space usage is only two words.
One way to leverage previous sketching research is to reduce complex sketching tasks to simpler ones, which are both well studied and widely deployed. We effect these reductions through a number of powerful emulation theorems, which create sketches for complex estimation tasks whose distribution is identical to an existing sketch.
Theorem 3 ($F_{\alpha}$-stable emulation, Section˜3.3).
Let $f$ be the characteristic exponent of any $d$-dimensional symmetric $\alpha$-stable process $X$. When $\alpha=2$, $f(x)=e^{-\frac{1}{2}t\langle x,Ax\rangle}$, where $A$ is a covariance matrix, and when $\alpha\in(0,2)$, $f(x)=\int_{\mathbb{S}_{n-1}}|\langle x,\xi\rangle|^{\alpha}\,\mu(d\xi)$, where $\mu$ is a symmetric, positive measure on $\mathbb{S}_{n-1}$. [^7] For the case $\alpha\in(0,2)$, only symmetric processes are considered here for simplicity. See [50] for the full characterization of stable processes. Let the Lévy-Stable sketch be parameterized by Lévy process $X$ on $\mathbb{R}^{d}$. Given any input vector $\mathbf{x}\in(\mathbb{R}^{d})^{n}$ and $\mathbf{x}^{\prime}\in\mathbb{R}^{n}$ such that $f(\mathbf{x})=\sum_{v\in[n]}|\mathbf{x}^{\prime}(v)|^{\alpha}$, Lévy-Stable with input $\mathbf{x}$ and Indyk’s $F_{\alpha}$-stable sketch with input $\mathbf{x}^{\prime}$ distribute identically.
Theorem˜3 serves as another illuminating example showing how the Lévy-Khintchine theorem helps to understand streaming sketching. Previously, only two classes of stable moments are considered: one dimensional $F_{\alpha}$-moments by Indyk [29], and multidimensional $F_{p,q}$-moments by Ganguly et al. [28] (see also [34]), which are sketched by algorithmic tricks of combining stable random variables. The Lévy-Stable sketch extends such tricks to all stable processes in a systematic way. For example, we can now estimate, using any estimator of Indyk’s sketch [29, 39], the $f$-moment of an $\mathbb{R}^{3}$-turnstile stream, where $f$ is the 1-stable function
$f(x)=\int_{\mathbb{S}_{2}}\frac{|\langle\xi,x\rangle|}{|\xi_{1}|^{2}+|\xi_{2}|+|\xi_{3}|^{1/2}}\,d\xi,$
and $\mathbb{S}_{2}$ is the unit $\mathbb{R}^{3}$-sphere. (It is 1-stable due to the numerator $|\langle\xi,x\rangle|$. The denominator can be any symmetric function of $\xi$ that is bounded away from 0 on the sphere.) We are not aware of any prior work that proved such stable $f$-moments could be sketched efficiently.
For the following two emulation theorems, let $G(z)=c+\gamma_{0}z+\int_{0}^{\infty}(1-e^{-zs})\,\nu(ds)$ be the Laplace exponent of any subordinator $X$ and parameterize the Lévy-based sketches by $X$. A cardinality sketch is Poissonized if each actual insertion is simulated by $\mathrm{Poisson}(1)$ insertions.
Theorem 4 (PCSA emulation, Section˜5).
Let vectors $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}_{+}^{n}$ be such that $G(\mathbf{x})=\|\mathbf{x}^{\prime}\|_{0}$. Then LévyPCSA (parameterized by $G$) on input $\mathbf{x}$ and Poissonized PCSA with input $\mathbf{x}^{\prime}$ distribute identically.
Theorem 5 (HyperLogLog emulation, Section˜5).
Let vectors $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}_{+}^{n}$ be such that $G(\mathbf{x})=\|\mathbf{x}^{\prime}\|_{0}$. Then LévyHyperLogLog (parameterized by $G$) with input $\mathbf{x}$ and Poissonized HyperLogLog with input $\mathbf{x}^{\prime}$ distribute identically.
The significance of the three emulation theorems is that, since the final distribution of the sketches are the same as the classic sketches, every analysis, estimator, or practical optimization developed over the years can be applied to Lévy-Stable, LévyPCSA, and LévyHyperLogLog for free, such as Li’s estimators for $F_{\alpha}$-stable sketches [39], the near-optimal $\tau$-GRA estimators [55] for PCSA and HyperLogLog, and maximum likelihood estimation for PCSA [47, 38]. In particular, the Fishmonger sketch [47] (and entropy-compressed version of PCSA with maximum likelihood estimation) can now estimate $G(\mathbf{x})$ with relative variance $1/m$ using $(1+o(1))m(H_{0}/I_{0})\approx 1.98m$ bits, for $m=\Omega(\log^{2}\log n)$. [^8] By definition $H_{0}=\frac{1}{\log 2}+\sum_{k=1}^{\infty}\frac{1}{k}\log_{2}\left(1+1/k\right)$ and $I_{0}=\pi^{2}/6$.
1.6 Organization
We first review Lévy processes and the Lévy-Khintchine representation theorem in Section˜2. We prove the two main theorems in Section˜3 and Section˜4 that connect the Lévy-Khintchine theorem to streaming sketches. Specifically, in Section˜3 we construct the Lévy-Tower sketch, which transforms any generic Lévy process $X$ into an $f_{X}$-moment sketch, where $f_{X}$ is the characteristic exponent of $X$. A specialization of Lévy-Tower is presented, call Lévy-Stable, which is more space efficient and applies whenever $X$ is a (multidimensional) stable process. In Section˜4 we construct the Lévy-Min-Sampler, which can be parameterized to sample elements according to any weight function $G$, where $G$ is the Laplace exponent of a subordinator.
In Section˜5, we use the $G$-transformation technique to reduce the problem of $G$-moment estimation to cardinality estimation; this leads to the LévyPCSA, LévyHyperLogLog, and Stable-HyperLogLog sketches. In Section˜6, we discuss the connection between previous sketches and Lévy processes in greater detail. In Section˜7, we discuss the problem of characterizing the set of tractable functions, and introduce the powerful Fourier-Hahn-Lévy method for expanding the range of the Lévy-Tower beyond Lévy-Khintchine-representable functions. We conclude in Section˜8 with some conjectures on the relationship between moment estimation, sampling, and Lévy processes.
2 Preliminaries: Lévy Processes and the Lévy-Khintchine Theorem
2.1 Lévy Processes on $\mathbb{R}^{d}$
We use Sato’s text [50] as our reference for the theory of Lévy processes.
Definition 3 (Lévy processes [50, page 3]).
A random process $X=(X_{t})_{t\geq 0}$ on $\mathbb{R}^{d}$ is a Lévy process if it has the following three properties.
(A) Stationary Increments.
$X_{t+s}-X_{t}\sim X_{s}$ for all $t,s\in\mathbb{R}_{+}$;
(B) Independent Increments.
for $0\leq t_{1}<t_{2}\ldots<t_{k}$, $X_{t_{1}},X_{t_{2}}-X_{t_{1}},\ldots,X_{t_{k}}-X_{t_{k-1}}$ are mutually independent;
(C) Stochastic Continuity.
$X_{0}=0$ almost surely and $\lim_{t\searrow 0}\mathbb{P}(|X_{t}|>\epsilon)=0$ for any $\epsilon>0$.
A process that satisfies (A) and (B) is called memoryless, i.e., conditioned on $X_{t^{*}}$, $(X_{t})_{t>t^{*}}$ is independent of $(X_{t})_{t<t^{*}}$. The primary way to study Lévy processes is through their characteristic functions. Refer to Table˜1 for notation.
Theorem 6 (Lévy-Khintchine representation [50, page 37]).
Any Lévy process $X=(X_{t})_{t\geq 0}$ on $\mathbb{R}^{d}$ can be identified by a triplet $(A,\nu,\gamma)$ where $A$ is a covariance matrix, $\gamma\in\mathbb{R}^{d}$, and $\nu$ is a measure on $\mathbb{R}$ such that
$\nu(\{0\})=0\quad\text{and}\quad\int_{\mathbb{R}^{d}}\min\{|s|^{2},1\}\,\nu(d{s})<\infty.$ (2)
The identification is through the characteristic function. For any $t\geq 0$ and $z\in\mathbb{R}^{d}$,
$\mathbb{E}e^{i\left\langle X_{t},z\right\rangle}$ $=\exp\left(-t\left(\frac{1}{2}\left\langle z,Az\right\rangle-i\left\langle\gamma,z\right\rangle+\int_{\mathbb{R}^{d}}(1+i\left\langle z,s\right\rangle\mathbbm{1}\!\left\{{|s|<1}\right\}-e^{i\left\langle z,s\right\rangle})\,\nu(ds)\right)\right).$ (3)
Conversely, any triplet $(A,\nu,\gamma)$ where $A$ is a covariance matrix, $\nu$ satisfies (2), and $\gamma\in\mathbb{R}^{d}$ corresponds to a Lévy process satisfying (3).
We call the exponent in (3) the characteristic exponent, denoted by $f_{X}(z)=-\log\mathbb{E}e^{i\left\langle X_{1},z\right\rangle}$.
Remark 1.
In the one-dimensional case, the matrix $A$ can be identified by the usual variance $\sigma^{2}$. Then we have for $z\in\mathbb{R}$,
$\mathbb{E}e^{izX_{t}}$ $=\exp\left(-t\left(\frac{1}{2}\sigma^{2}z^{2}-i\gamma z+\int_{\mathbb{R}}(1+izs\mathbbm{1}\!\left\{{|s|<1}\right\}-e^{izs})\,\nu(ds)\right)\right).$ (4)
Moreover, the characteristic exponent $f_{X}(z)$ is real if and only if the Lévy measure $\nu$ is symmetric around 0 [50], in which case the characteristic exponent can be expressed as
$\mathbb{E}e^{izX_{t}}$ $=\exp\left(-t\left(Ax^{2}+2\int_{0}^{\infty}(1-\cos(sz))\nu(ds)\right)\right)$ (5)
for some constant $A$ and Lévy measure $\nu$.
We note some properties of the characteristic exponent $f$.
Lemma 1.
Let $X$ be any Lévy process on $\mathbb{R}^{d}$ and $f$ be its characteristic exponent.
  • $\mathfrak{Re}f\geq 0$. $\mathfrak{Re}f=0$ if and only if $X$ is a deterministic drift, i.e., $X_{t}=\gamma t$ for all $t\geq 0$.
  • For any $x\in\mathbb{R}^{d}$, $f(-x)=\overline{f(x)}$, the complex conjugate of $f(x)$.
We list some common one-dimensional Lévy processes in Table˜2 that will be frequently used later.
Process Characteristic Exponent Notes
Linear drift $i\gamma z$ $\gamma\in\mathbb{R}_{+}$
$\alpha$-stable $|z|^{\alpha}$ $\alpha\in(0,2]$
Poisson $1-e^{iz}$
Compound Poisson $\int_{-\infty}^{\infty}(1-e^{izs})\,\nu(ds)$ $\nu(\mathbb{R})<\infty$ is the jump rate,
$\nu(\mathbb{R})^{-1}\nu$ is the jump distribution
Symmetric Compound Poisson $2\int_{0}^{\infty}(1-\cos(zs))\,\nu(ds)$ $\nu(\mathbb{R})<\infty$
Table 2: Common one-dimensional Lévy processes.
2.2 Subordinators: Lévy Processes on $\mathbb{R}_{+}\cup\{\infty\}$
One-dimensional, non-negative Lévy processes occupy a special place in the theory of Lévy processes [50]. They are also known as subordinators for reasons that will be explained shortly. As all subordinators are supported on $\mathbb{R}_{+}\cup\{\infty\}$ they can be characterized via their Laplace transforms.
Theorem 7 (Lévy-Khintchine for subordinators [50, page 138]).
Any subordinator $X=(X_{t})_{t\geq 0}$ on $\mathbb{R}_{+}\cup\{\infty\}$ can be identified by a triplet $(c,\nu,\gamma_{0})$ where $c,\gamma_{0}\geq 0$, and $\int_{0}^{\infty}\min\{s,1\}\,\nu(ds)<\infty$, such that for any $t,z\in\mathbb{R}_{+}$,
$\mathbb{E}e^{-zX_{t}}$ $=\exp\left(-t\left(c\mathbbm{1}\!\left\{{z>0}\right\}+\gamma_{0}z+\int_{0}^{\infty}(1-e^{-zs})\,\nu(ds)\right)\right).$ (6)
Conversely, given any $c,\gamma_{0}\geq 0$, and measure $\nu$ such that $\int_{0}^{\infty}\min\{s,1\}\,\nu(ds)<\infty$, there is a corresponding subordinator satisfying (6).
The exponent in (6) is called the Laplace exponent, denoted as $G_{X}(z)=-\log\mathbb{E}e^{-zX_{1}}$. Subordinators are necessarily non-decreasing. The parameter $c$ is called the kill rate, and only comes into play when it is possible for $X_{t}$ to reach $\infty$, effectively killing the process. Sato’s text [50] only considers processes with $c=0$. However, we do want to consider killed processes (those with $c>0$) here because, as we will see later, popular sketches like PCSA and HyperLogLog are induced by killed processes. There is a standard way to kill a process.
Lemma 2 (How to kill a Lévy process; see [57]).
We are given a subordinator $X$ with Laplace exponent $G:\mathbb{R}_{+}\to\mathbb{R}_{+}$. Fix any $c>0$, let $Y$ be an independent $\mathrm{Exp}(c)$ random variable, and define a new process $(X^{\prime}_{t})_{t\geq 0}$ on $\mathbb{R}_{+}\cup\{\infty\}$ as
$X^{\prime}_{t}=\begin{cases}X_{t},&t<Y,\\ \infty,&t\geq Y.\end{cases}$
Then for $z\in\mathbb{R}$,
$\mathbb{E}e^{-zX^{\prime}_{t}}=e^{-t(G(z)+c\mathbbm{1}\!\left\{{z>0}\right\})},$
where $e^{-z\infty}=\mathbbm{1}\!\left\{{z=0}\right\}$.
A non-negative Lévy process is non-decreasing and can therefore be used to index time within another Lévy process. This is called subordination. For example, if $(X_{t})_{t\geq 0}$ is a Wiener process/Brownian motion and $(Z_{t})_{t\geq 0}$ is a Poisson process with rate 1, then $(X_{Z_{t}})_{t\geq 0}$ ($X$ subordinated by $Z$) is piecewise constant. It increments the “clock” in the Wiener process by 1 at intervals that are independent $\mathrm{Exp}(1)$ random variables.
Theorem 8 (Subordination [50, page 198]).
Let $X=(X_{t})$ be a Lévy process on $\mathbb{R}^{d}$ and $Z=(Z_{t})_{t\geq 0}$ be a subordinator on $\mathbb{R}_{+}$, with $c=0$. Then $(X_{Z_{t}})_{t\geq 0}$ is a Lévy process on $\mathbb{R}^{d}$ such that for any $t\geq 0$ and $z\in\mathbb{R}^{d}$,
$\mathbb{E}e^{i\left\langle X_{Z_{t}},z\right\rangle}$ $=\exp(-tG_{Z}(f_{X}(z))),$
where $f_{X}$ is the characteristic exponent of $X$ and $G_{Z}$ is the Laplace exponent of $Z$.
3 Lévy-Tower and Lévy-Stable
3.1 Lévy-Tower Sketches
We now present a sketch that is induced by a generic Lévy process on $\mathbb{R}^{d}$, which consists of two parameters.
  • An accuracy parameter $m\in\mathbb{Z}_{+}$ which corresponds to the number of subsketches in classic settings.
  • A Lévy process $X=(X_{t})_{t\geq 0}$ on $\mathbb{R}^{d}$ with characteristic exponent $f(z)=-\log\mathbb{E}e^{i\left\langle X_{1},z\right\rangle}$.
Recall that $\mathbb{T}$ is the complex unit circle.
Definition 4 ($(f,m)$-Lévy-Tower).
Let $f$ be the characteristic exponent of a Lévy process $X$ on $\mathbb{R}^{d}$ and $m\in\mathbb{Z}_{+}$. An abstract $(f,m)$-Lévy-Tower is an infinite vector $S=(S_{k}^{(j)})_{k\in\mathbb{Z},j\in[m]}\subset\mathbb{T}^{\mathbb{Z}}$, initialized as all zero. For any element $v\in[n]$ and $y\in\mathbb{R}^{d}$, a vector update $\mathbf{x}(v)\leftarrow\mathbf{x}(v)+y$ is effected by:
$\mathsf{Update}(v,y):$
For each $k\in\mathbb{Z}$ and $j\in[m]$, $S_{k}^{(j)}\leftarrow S_{k}^{(j)}+\left\langle y,X_{2^{-k}}^{(v,j)}\right\rangle\mod 2\pi$, where $X^{(v,j)}=(X^{(v,j)}_{t})_{t\geq 0}$ is an i.i.d. copy of the Lévy process $X$ with characteristic exponent $f$.
Remark 2.
If $|f(\mathbf{x})|$ is in the range $1,\operatorname{poly}(n)]$, then it suffices to store levels $k\in[0,O(\log n)]$. Therefore a Lévy-Tower takes $O(m\log n)$ words of space and each word stores a number in $0,2\pi)$.
Many sketches follow the three-step design template: subsample, randomly project, and sum. The Lévy-Tower can be regarded as doing the subsample and projection steps at the same time, where the measurement time $2^{-k}$ is analogous to a subsampling probability.
We characterize the distribution of the sketch as follows.
Figure 1: Lévy-Tower with $m=3$ and $\mathbf{x}=(1,0,\ldots,0)$. From left to right: linear drift, the 1-stable Cauchy process, the 2-stable Wiener process/Brownian motion, the Poisson process with rate 1, and the Poisson process with rate 2. Different Lévy processes have different “sensitivities” for a target function-moment. For example, linear drift is only sensitive to the sum of the vector and insensitive to how the values are distributed. The Cauchy process is only sensitive to the $F_{1}$-moment, while the Wiener process/Brownian motion is only sensitive to the $F_{2}$-moment. Poisson processes are sensitive to the support size of $\mathbf{x}$ and at the same time leak information about other $f$-moments.
Lemma 3.
Fix a vector $\mathbf{x}=(\mathbf{x}(v))_{v\in[n]}\in(\mathbb{R}^{d})^{n}$. For any $k\in\mathbb{Z}$, we have
$S_{k}$ $\sim\sum_{v\in[n]}\left\langle X_{2^{-k}}^{(v)},\mathbf{x}(v)\right\rangle$
$\mathbb{E}e^{iS_{k}}$ $=e^{-2^{-k}f(\mathbf{x})}.$
Proof.
The first statement is trivially true since the sketch is linear. For the second, note that the $X^{(v)}$ are i.i.d., and we have
$\mathbb{E}e^{iS_{k}}$ $=\mathbb{E}e^{i\sum_{v\in[n]}\left\langle X_{2^{-k}}^{(v)},\mathbf{x}(v)\right\rangle}$
$=\prod_{v\in[n]}\mathbb{E}e^{i\left\langle X_{2^{-k}}^{(v)},\mathbf{x}(v)\right\rangle}$ (by independence)
$=\prod_{v\in[n]}e^{-2^{-k}f(\mathbf{x}(v))}$ (by Lévy-Khintchine)
$=e^{-2^{-k}f(\mathbf{x})}.$ (definition of $f(\mathbf{x})$)
Note that the $f$-moment lies exactly in the exponent of $\mathbb{E}e^{iS_{k}}$. We now formally present a method that can recover an accurate estimate of the $f$-moment from $S_{k}$, or more accurately, $m$ i.i.d. copies $S_{k}^{(1)},\ldots,S_{k}^{(m)}$.
3.2 Estimation of Lévy-Tower
Recall from Section˜1.3 that $C_{t}=\sum_{v\in[n]}\left\langle\mathbf{x}(v),X_{t}^{(v)}\right\rangle$ is the linear projection measured at time $t$. The estimator must pick a suitable $t$ and infer an estimate of the $f$-moment from $C_{t}$. In the Lévy-Tower, we store $S_{k}=C_{2^{-k}}$ for all $k\in\mathbb{Z}$. We first prove the concentration of the empirical mean at each level.
Lemma 4.
For any $t\geq 0$ and $m$ i.i.d. copies $C_{t}^{(1)},\ldots,C_{t}^{(m)}$,
$\mathbb{P}\left(\left|\frac{1}{m}\sum_{j=1}^{m}e^{iC_{t}^{(j)}}-e^{-tf(\mathbf{x})}\right|>\eta\right)\leq\frac{2t|f(\mathbf{x})|}{m\eta^{2}}.$
Proof.
By Chebyshev’s inequality, we have
$\mathbb{P}\left(\left|\frac{1}{m}\sum_{j=1}^{m}e^{iC_{t}^{(j)}}-e^{-tf(\mathbf{x})}\right|>\eta\right)\leq\frac{\mathbb{V}e^{iC_{t}}}{m\eta^{2}}.$
Furthermore, since $\mathbb{E}e^{iC_{t}}=e^{-tf(\mathbf{x})}$, $|e^{iC_{t}}|=1$, and $|\mathbb{E}e^{iC_{t}}|^{2}=e^{-2t\cdot\mathfrak{Re}f(\mathbf{x})}$, we have
$\mathbb{V}e^{iC_{t}}$ $=\mathbb{E}\left|e^{iC_{t}}\right|^{2}-\left|\mathbb{E}e^{iC_{t}}\right|^{2}=1-e^{-2t\cdot\mathfrak{Re}f(\mathbf{x})}\leq 2t\cdot\mathfrak{Re}f(\mathbf{x})\leq 2t|f(\mathbf{x})|.$
Lemma 5.
With probability at least $99/100$, for any $k\geq\log_{2}|f(\mathbf{x})|$, we have
$\left|\frac{1}{m}\sum_{j=1}^{m}e^{iS_{k}^{(j)}}-e^{-2^{-k}f(\mathbf{x})}\right|\leq O(1/\sqrt{m}).$
Proof.
By the union bound and Lemma˜4,
$\mathbb{P}\left(\exists k\geq\log_{2}|f(\mathbf{x})|.\;\left|\frac{1}{m}\sum_{j=1}^{m}e^{iS_{k}^{(j)}}-e^{-2^{-k}f(\mathbf{x})}\right|>\eta\right)$
$\leq\sum_{k=\log_{2}|f(\mathbf{x})|}^{\infty}\frac{2\cdot 2^{-k}|f(\mathbf{x})|}{m\eta^{2}}$
$=O\left(\frac{1}{m\eta^{2}}\right).$
Thus, it suffices to choose $\eta=O(1/\sqrt{m})$ for the probability above to be at most $1/100$. ∎
We consider the complex logarithm defined on $\mathbb{C}$ except for the negative real line.
Lemma 6.
Let $z\in\mathbb{C}$ with $|z|\leq 1$ and $\mathfrak{Re}(z)\geq 0$, then $|1-e^{-z}|\in(1/2,1]\cdot|z|$.
Lemma 7.
Let $x,y\in\{z\in\mathbb{C}:|1-z|\leq 0.98\}$, then $|\log x-\log y|<50|x-y|$.
Input: A Lévy-Tower $(S_{k}^{(j)})_{k\in[K],j\in[m]}$ sketch of $\mathbf{x}\in(\mathbb{R}^{d})^{n}$, parameterized by Lévy process $X$ with characteristic exponent $f=f_{X}$.
Output: An estimate $\widehat{f(\mathbf{x})}$ such that if $\log_{2}|f(\mathbf{x})|<K$ and $m=\Omega(1)$, then with probability 99/100, $|\widehat{f(\mathbf{x})}-f(\mathbf{x})|=O(|f(\mathbf{x})|/\sqrt{m})$.
for $W$ from $K$ down to 1 do
 $Y_{W}=\frac{1}{m}\sum_{j=1}^{m}e^{iS_{W}^{(j)}}$
 if $|1-Y_{W}|\geq 0.12$ then
    return $\widehat{f(\mathbf{x})}=-2^{W}\log Y_{W}$
   end if
end for
return $2\log Y_{1}$
// fail outcome
Algorithm 1 ff-Moment-Estimation
Theorem 9.
Consider a Lévy-Tower sketch $(S_{k}^{(j)})$ of a vector $\mathbf{x}\in(\mathbb{R}^{d})^{n}$, where $k\in[K],j\in[m]$, and $K>\log_{2}|f(\mathbf{x})|$. With probability at least 99/100, $f$-Moment-Estimation (Algorithm˜1) returns an estimate $\widehat{f(\mathbf{x})}$ such that $|\widehat{f(\mathbf{x})}-f(\mathbf{x})|\leq O(|f(\mathbf{x})|/\sqrt{m})$.
Proof.
We analyze the behavior of the estimator conditioned on event $\mathcal{E}$ holding. Lemma˜5 states that $\mathbb{P}(\mathcal{E})\geq 99/100$.
 : $\text{$\mathcal{E}$ : For all $k\in[\log_{2}|f(\mathbf{x})|,K]$,}\quad\left|Y_{k}-e^{-2^{-k}f(\mathbf{x})}\right|\leq O(1/\sqrt{m}).$ (7)
For $k\in[\log_{2}|f(\mathbf{x})|,K]$, we have $|2^{-k}f(\mathbf{x})|\leq 1$ and by Lemma˜1 we know $\mathfrak{Re}(f(\mathbf{x}))\geq 0$. Thus we can bound the distance from $Y_{k}$ to 1 both from below and above, as follows.
$|1-Y_{k}|$ $\in\left|1-e^{-2^{-k}f(\mathbf{x})}\right|\pm\left|e^{-2^{-k}f(\mathbf{x})}-Y_{k}\right|$ triangle inequality
$\subset\left|1-e^{-2^{-k}f(\mathbf{x})}\right|\pm O(1/\sqrt{m})$
$\subset(1/2,1]\left|2^{-k}f(\mathbf{x})\right|\pm O(1/\sqrt{m}).$ (8)
Let $m$ be large enough such that the $O(1/\sqrt{m})$ term is at most $0.1$. When the estimate is returned, $W$ is the largest index such that $|1-Y_{W}|\geq 0.12$. It follows that $W\geq\lceil\log_{2}|f(\mathbf{x})|\rceil$ since by Eq.˜8,
$|1-Y_{\lceil\log_{2}|f(\mathbf{x})|\rceil}|\in(1/2,1]\left|2^{-{\lceil\log_{2}|f(\mathbf{x})|\rceil}}f(\mathbf{x})\right|\pm O(1/\sqrt{m})>1/4-0.1=0.15,$
implying the algorithm halted with $W\geq\lceil\log_{2}|f(\mathbf{x})|\rceil$. We further claim that $W-\lceil\log_{2}|f(\mathbf{x})|\rceil=O(1)$ and specifically that
$2^{-W}|f(\mathbf{x})|$ $\in[0.02,0.88].$ (9)
If it were the case that $2^{-W}|f(\mathbf{x})|<0.02$ then Eq.˜8 implies that
$|1-Y_{W}|<0.02+0.1=0.12,$
which contradicts the choice of $W$. On the other hand, $2^{-W}|f(\mathbf{x})|$ cannot be bigger than 0.88 for otherwise one would have $2^{-(W+1)}|f(\mathbf{x})|\geq 0.44$ and Eq.˜8 implies
$|1-Y_{W+1}|>\frac{1}{2}\left|2^{-(W+1)}f(\mathbf{x})\right|>0.22-0.1=0.12,$
also contradicting the choice of $W$. By Lemma˜6 and Eq.˜9 we have
$\left|1-e^{-2^{-W}f(\mathbf{x})}\right|\leq\left|2^{-W}f(\mathbf{x})\right|\leq 0.88,$
and by Eq.˜8 we have $|1-Y_{W}|\leq 0.88+0.1=0.98$. Applying Lemmas˜5 and 7, we have
$\left|\log Y_{W}-\left(-2^{-W}f(\mathbf{x})\right)\right|$ $\leq 50\left|Y_{W}-e^{-2^{-W}f(\mathbf{x})}\right|$ Lemma˜7, $x=Y_{W},y=e^{-2^{-W}f(\mathbf{x})}$,
$=O(1/\sqrt{m})$ Event $\mathcal{E}$ (Lemma˜5).
Multiplying by $2^{W}=\Theta(|f(\mathbf{x})|)$ we conclude that $\widehat{f(\mathbf{x})}$ achieves the desired approximation, conditioned on event $\mathcal{E}$.
$\left|\widehat{f(\mathbf{x})}-f(\mathbf{x})\right|=\left|-2^{W}\log Y_{W}-f(\mathbf{x})\right|\leq O(2^{W}/\sqrt{m})=O(|f(\mathbf{x})|/\sqrt{m}).$
3.3 Lévy-Stable Sketches
Recall that a $d$-dimensional Lévy process $X$ is $\alpha$-stable if for any $t\geq 0$, $X_{t}\sim t^{1/\alpha}X_{1}$. We call the characteristic exponents of stable processes stable moments. Stable moments are of special interest in the context of streaming sketches since there is no need to store the whole tower; only $m$ i.i.d. samples suffice to return an estimate with relative error $O(1/\sqrt{m})$. [^9] The Lévy-Towers induced by stable processes are self-similar: all registers in the tower are identically distributed with a proper normalization. The only symmetric one-dimensional stable processes are $\alpha$-stable random processes, which correspond to Indyk’s sketches [29]. On the other hand, there is a rich class of higher dimensional stable processes, the one implicit in Ganguly et al. [28] for estimating $F_{p,q}$ hybrid moments being just a single special case.
Theorem 10 (Lévy-Khintchine for symmetric stable processes [50, page 86]).
Let $X$ be a Lévy process on $\mathbb{R}^{d}$ and $x\in\mathbb{R}^{d}$. $X$ is $2$-stable if and only if $\mathbb{E}e^{i\langle x,X_{t}\rangle}=e^{-\frac{1}{2}t\langle x,Ax\rangle}$ for some covariance matrix $A$. If $X$ is symmetric, then it is $\alpha$-stable for $\alpha\in(0,2)$ if and only if there is a finite, positive, symmetric measure $\mu$ on the sphere $\mathbb{S}_{d-1}=\{x\in\mathbb{R}^{d}:|x|=1\}$ such that
$\mathbb{E}e^{i\langle x,X_{t}\rangle}=\exp\left(-t\int_{\mathbb{S}_{d-1}}|\langle x,\xi\rangle|^{\alpha}\,\mu(d\xi)\right).$
Remark 3.
For simplicity we only consider symmetric processes here when $\alpha\in(0,2)$ so that the characteristic exponent $f_{X}(x_{1},\ldots,x_{d})=f_{X}(|x_{1}|,\ldots,|x_{d}|)$ for any $x\in\mathbb{R}^{d}$. See [50] for the full characterization of (not necessarily symmetric) stable processes over $\mathbb{R}^{d}$.
It turns out that any multidimensional $\alpha$-stable moment is as simple to estimate as the one-dimensional $F_{\alpha}$-moment, since it is possible to maintain a random variable distributed as $f(\mathbf{x})^{1/\alpha}Y_{\alpha}$ where $Y_{\alpha}$ is a unit $\alpha$-stable random variable. We now give the formal definition of the Lévy-Stable sketch together with the proof of the emulation theorem.
Definition 5 ($(f,m)$-Lévy-Stable).
Let $f$ be the characteristic exponent of an $\alpha$-stable Lévy process $X$ on $\mathbb{R}^{d}$ and $m\in\mathbb{Z}_{+}$. An $(f,m)$-Lévy-Stable sketch is a vector of $m$ registers $T=(T_{1},\ldots,T_{m})\in\mathbb{R}^{m}$, initialized as all zero. For any element $v\in[n]$ and $y\in\mathbb{R}^{d}$, the vector update $\mathbf{x}(v)\leftarrow\mathbf{x}(v)+y$ is effected by:
$\mathsf{Update}(v,y):$
For each $k\in[m]$, $T_{k}\leftarrow T_{k}+\left\langle y,X_{k}^{(v)}-X_{k-1}^{(v)}\right\rangle$, where $X^{(v)}=(X^{(v)}_{t})_{t\geq 0}$ is an independent copy of the Lévy process $X$ with characteristic exponent $f$, indexed by $v$.
Proof of Theorem˜3.
Since $X$ is Lévy, $X_{1},X_{2}-X_{1},\ldots,X_{m}-X_{m-1}$ are i.i.d. with the same distribution as $X_{1}$, so it suffices to look at the first register $T_{1}=\sum_{v\in[n]}\left\langle\mathbf{x}(v),X^{(v)}_{1}\right\rangle$. Recall that for a stream with frequency vector $\mathbf{x}^{\prime}$, Indyk [29] stores $I_{1}=\sum_{v\in[n]}\mathbf{x}^{\prime}(v)Y^{(v)}_{\alpha}$ where $Y_{\alpha}$ is a unit $\alpha$-stable random variable. By the properties of stable random variables, we have
$I_{1}=\sum_{v\in[n]}\mathbf{x}^{\prime}(v)Y^{(v)}_{\alpha}\sim\left(\sum_{v\in[n]}|\mathbf{x}^{\prime}(v)|^{\alpha}\right)^{1/\alpha}Y_{\alpha}$
To show the simulation theorem, we need to prove
$T_{1}=\sum_{v\in[n]}\left\langle\mathbf{x}(v),X^{(v)}_{1}\right\rangle\sim\left(\sum_{v\in[n]}f(\mathbf{x}(v))\right)^{1/\alpha}Y_{\alpha},$
so that $T_{1}\sim I_{1}$ if the $f$-moment of $\mathbf{x}$ is equal to the $F_{\alpha}$-moment $\mathbf{x}^{\prime}$. We compute the characteristic function
$\mathbb{E}e^{izT_{1}}$ $=\mathbb{E}e^{iz\sum_{v\in[n]}\left\langle\mathbf{x}(v),X^{(v)}_{1}\right\rangle}=\prod_{v\in[n]}\mathbb{E}e^{i\left\langle\mathbf{x}(v),zX^{(v)}_{1}\right\rangle}$ (by independence)
Since $X^{(v)}$ is $\alpha$-stable, we have $zX_{1}^{(v)}\sim X_{z^{\alpha}}^{(v)}$
$=\prod_{v\in[n]}\mathbb{E}e^{i\left\langle\mathbf{x}(v),X^{(v)}_{z^{\alpha}}\right\rangle}$
$=\prod_{v\in[n]}e^{-z^{\alpha}f(\mathbf{x}(v))}$ (by Lévy-Khintchine)
$=e^{-z^{\alpha}f(\mathbf{x})}=e^{-(zf(\mathbf{x})^{1/\alpha})^{\alpha}}$
$=\mathbb{E}e^{izf(\mathbf{x})^{1/\alpha}Y_{\alpha}}.$ (by Lévy-Khintchine)
Thus $T_{1}$ has the same characteristic function with $f(\mathbf{x})^{1/\alpha}Y_{\alpha}$, which implies $T_{1}\sim f(\mathbf{x})^{1/\alpha}Y_{\alpha}$. ∎
Remark 4.
We remark here on the differences between the Lévy-Tower sketch and the Lévy-Stable sketch.
  • Lévy-Tower estimates a generic function moment $f_{X}$ where $X$ is the characteristic exponent of any Lévy process. Lévy-Tower stores projections at multiple times of the process and each projection can be taken modulo $2\pi$.
  • Lévy-Stable estimates a function moment $f_{X}$ where $X$ is the characteristic exponent of a stable Lévy process. Lévy-Stable stores projections at unit time of the process and each projection is in $\mathbb{R}$. Lévy-Stable generalizes the stable sketches of Indyk [29] and Ganguly, Bansal, and Dube [28].
4 The Lévy-Min-Sampler
The main takeaway message from Section˜3 is that in the $\mathbb{R}^{d}$-turnstile model, there is a uniform way to solve $f$-moment estimation whenever $f$ is the characteristic exponent of a Lévy process. The thesis of this section is that non-negative one-dimensional Lévy processes (subordinators) play a similar role in $G$-sampling and $G$-moment estimation in incremental streams.
4.1 The Raw Lévy-Min-Sampler
Recall that a generic min sketch (Definition˜2) is parameterized by a hash function $H:[n]\times\mathbb{R}_{+}\times\mathbb{Z}_{+}\to\mathbb{R}_{+}$. The sketch stores a pair $(v_{*},h_{*})$ initialized as $(\perp,\infty)$ and then reads the stream input at time $k=1,2,3,\ldots$
$\mathsf{Update}(v,\Delta)$ (issued at time $k$):
if $H(v,\Delta,k)<h_{*}$, then set $(v_{*},h_{*})\leftarrow(v,H(v,\Delta,k))$.
Based on the discussion in Section˜1.4, we now consider hash functions induced by subordinators.
Theorem 11 (Raw Lévy-Min-Sampler).
For $u\in[n],\Delta\in\mathbb{R}_{+}$, and $k\in\mathbb{Z}_{+}$, define
$H(u,\Delta,k)=\inf\left\{t\geq 0:X^{(u)}_{t}>Z_{k}/\Delta\right\},$
where the $X^{(u)}=(X^{(u)}_{t})_{t\geq 0}$ are i.i.d. subordinators with Laplace exponent $G:\mathbb{R}_{+}\to\mathbb{R}_{+}$ and the $Z_{k}\sim\mathrm{Exp}(1)$ are i.i.d. standard exponential random variables. Then the min sketch induced by $H$ is a $G$-sampler. I.e., the resulting state $(v_{*},h_{*})$ satisfies the following.
  • For any $u\in[n]$, $\mathbb{P}(v_{*}=u)=G(\mathbf{x}(u))/G(\mathbf{x})$.
  • $h_{*}\sim\mathrm{Exp}(G(\mathbf{x}))$.
The proof of Theorem˜11 makes use of the following standard properties of exponential random variables.
Lemma 8 (Properties of exponential random variables).
Let $Y_{1},\ldots,Y_{k}$ be independent random variables where $Y_{j}\sim\mathrm{Exp}(\lambda_{j})$ for $j\in[k]$. Then
  1. 1.
    $\min\{Y_{1},\ldots,Y_{k}\}\sim\mathrm{Exp}(\lambda_{1}+\cdots+\lambda_{k})$.
  2. 2.
    $\mathbb{P}\left(Y_{u}<\min_{j\neq u}Y_{j}\right)=\frac{\lambda_{u}}{\sum_{j\in[k]}\lambda_{j}}$.
Proof of Theorem˜11.
Suppose element $u$ is updated at times $(k_{j})$ with increments $(\Delta_{j})$, where $\sum_{j}\Delta_{j}=\mathbf{x}(u)$. Observe that for any $z\in\mathbb{R}_{+}$,
$\mathbb{P}\left(\min_{j}\{H(u,\Delta_{j},k_{j})\}\geq z\right)$ $=\mathbb{P}\left(\inf\left\{t\geq 0:X^{(u)}_{t}>\min_{j}\{Z_{k_{j}}/\Delta_{j}\}\right\}\geq z\right)$
$=\mathbb{P}\left(X^{(u)}_{z}\leq\min_{j}\{Z_{k_{j}}/\Delta_{j}\}\right).$
Since the $Z_{k_{j}}$ are i.i.d. $\mathrm{Exp}(1)$ random variables, $\min_{j}\left\{Z_{k_{j}}/\Delta\right\}\sim\mathrm{Exp}\left(\sum_{j}\Delta_{j}\right)=\mathrm{Exp}(\mathbf{x}(u))$. Continuing,
$=\mathbb{P}(X^{(u)}_{z}\leq\mathrm{Exp}(\mathbf{x}(u)))$
$=\mathbb{E}\mathbb{P}\left({X^{(u)}_{z}\leq\mathrm{Exp}(\mathbf{x}(u))}\;\middle|\;X^{(u)}_{z}\right)$
$=\mathbb{E}e^{-\mathbf{x}(u)X^{(u)}_{z}}$,
and since $X^{(u)}$ has Laplace exponent $G$, this is equal to
$=e^{-zG(\mathbf{x}(u))}.$
By the CDF of the exponential distribution,
$\min_{j}\{H(u,\Delta_{j},k_{j})\}$ $\sim\mathrm{Exp}(G(\mathbf{x}(u))).$
4.2 Stable-Min-Sampler
When $G(x)=x^{\alpha}$, $\alpha\in(0,1)$, we can obtain a much simpler emulation. Recall that a standard one-sided $\alpha$-stable random variable $W$ is one for which $\mathbb{E}e^{-zW}=e^{-z^{\alpha}}$.
Theorem 12 (Stable-Min-Sampler).
Let $\alpha\in(0,1)$. For $u\in[n]$ and $k\in\mathbb{Z}_{+}$, define
$H(u,\Delta,k)=\left(\frac{Z_{k}}{\Delta W^{(u)}}\right)^{\alpha},$
where $W^{(u)}$s are i.i.d. standard one-sided $\alpha$-stable random variables and the $Z_{k}\sim\mathrm{Exp}(1)$ are i.i.d. standard exponential random variables. Then the min sketch induced by $H$ is an $F_{\alpha}$-sampler. I.e., the resulting state $(v_{*},h_{*})$ satisfies the following.
  • For any $u\in[n]$, $\mathbb{P}(v_{*}=u)=\mathbf{x}(u)^{\alpha}/\sum_{v\in[n]}\mathbf{x}(v)^{\alpha}$.
  • $h_{*}\sim\mathrm{Exp}\left(\sum_{v\in[n]}\mathbf{x}(v)^{\alpha}\right)$.
Proof.
Note that
$\mathbb{P}\left(\left(\frac{\mathrm{Exp}(\mathbf{x}(u))}{W^{(u)}}\right)^{\alpha}\geq z\right)$ $=\mathbb{P}\left(\mathrm{Exp}(\mathbf{x}(u))\geq z^{1/\alpha}W^{(u)}\right)$
$=\mathbb{E}\left(\mathbb{P}(\mathrm{Exp}(\mathbf{x}(u))\geq z^{1/\alpha}W^{(u)})\;\middle|\;W^{(u)}\right)$
$=\mathbb{E}e^{-z^{1/\alpha}\mathbf{x}(u)W^{(u)}}$
$=e^{-z\mathbf{x}(u)^{\alpha}}.$
4.3 Lévy-Min-Sampler
For a generic weight function $G$, the raw Lévy-Min-Sampler can be simplified since we only need $\min_{j}\{H(u,\Delta_{j},k_{j})\}$ to be distributed as $\mathrm{Exp}(G(\mathbf{x}(u)))$, which can be simulated through level functions. Level functions are similar to the score functions used by Cohen [15], which map the indexed hash values and fresh randomness into one number. We now construct level functions from subordinators.
Definition 6 (level function induced by $G$).
A level function $\ell$ is a map from $(0,\infty)\times(0,1)$ to $0,\infty]$ such that $\ell(x,y)$ is increasing in both $x$ and $y$. Given a subordinator $X=(X_{t})_{t\geq 0}$ with Laplace exponent $G:\mathbb{R}_{+}\to\mathbb{R}_{+}$, the induced level function is defined to be
$\ell_{G}(x,y)$ $=\inf\{z:\mathbb{P}(X_{z}\geq x)\geq y\},$
for any $x,y\in\mathbb{R}_{+}$.
The final Lévy-Min-Sampler using level functions is presented in Algorithm˜2 and Theorem˜13.
Theorem 13 (Lévy-Min-Sampler).
Let $H:[n]\to[0,1]$ be a uniformly random hash function. Let $G:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be the Laplace exponent of any subordinator. The Lévy-Min-Sampler is a pair $(v_{*},h_{*})$ initialized as $(\perp,\infty)$. A vector update $\mathbf{x}(v)\leftarrow\mathbf{x}(v)+\Delta$ at time $k$ is effected by:
$\mathsf{Update}(v,\Delta)$ (issued at time $k$):
If $\ell_{G}\left(\frac{Z_{k}}{\Delta},H(v)\right)<h_{*}$, then set $(v_{*},h_{*})\leftarrow\left(v,\ell_{G}\left(\frac{Z_{k}}{\Delta},H(v)\right)\right)$,
where the $Z_{k}\sim\mathrm{Exp}(1)$ are i.i.d. standard exponential random variables. When the input vector is $\mathbf{x}\in\mathbb{R}_{+}^{n}$, the resulting sampler state satisfies the following.
  • For any $u\in[n]$, $\mathbb{P}(v_{*}=u)=G(\mathbf{x}(u))/G(\mathbf{x})$.
  • $h_{*}\sim\mathrm{Exp}(G(\mathbf{x}))$.
Sketch : $(v_{*},h_{*})$, initialized as $(\perp,\infty)$
Hash function : $H:[n]\to\mathrm{Uniform}(0,1)$
Result: Sample an element $u$ with prob. $G(\mathbf{x}(u))/\sum_{v\in[n]}G(\mathbf{x}(v))$
// upon update $\mathbf{x}(v)\leftarrow\mathbf{x}(v)+\Delta$
Update $(v,\Delta)$
 $h\leftarrow\ell_{G}(\mathrm{Exp}(\Delta),H(v))$
 // $\mathrm{Exp}(\Delta)$ is freshly sampled
 if $h<h_{*}$ then
    $(v_{*},h_{*})\leftarrow(v,h)$
   end if
// upon sample
Sample ()(\,)
 return $v_{*}$
Algorithm 2 Lévy-Min-Sampler. $\ell_{G}$ is the level function for the weight function $G$ defined in Definition˜6.
Proof.
It suffices to prove that whenever $Z\sim\mathrm{Exp}(\lambda)$ and $Y\sim\mathrm{Uniform}(0,1)$, that $\ell_{G}(Z,Y)\sim\mathrm{Exp}(G(\lambda))$. Indeed, let $X=(X_{t})_{t\geq 0}$ be the subordinator with Laplace exponent $G$. We have, for any $w>0$,
$\mathbb{P}(\ell_{G}(Z,Y)\geq w)$ $=\mathbb{P}(\inf\{z:\mathbb{P}(X_{z}\geq Z)\geq Y\}\geq w)$ (by definition of $\ell_{G}$)
$=\mathbb{P}(\mathbb{P}(X_{w}\geq Z)<Y)$ ($(X_{t})$ non-decreasing)
$=\mathbb{P}(X_{w}<Z)$ ($Y\in\mathrm{Uniform}(0,1)$)
$=\mathbb{E}(\mathbb{P}(X_{w}<Z)\mid X_{w})$ (law of total probability)
$=\mathbb{E}(e^{-\lambda X_{w}}\mid X_{w})$
$=\mathbb{E}e^{-\lambda X_{w}}\>=\>e^{-wG(\lambda)}.$
By the CDF of the exponential distribution, $\ell_{G}(Z,Y)\sim\mathrm{Exp}(G(\lambda))$. ∎
We demonstrate by examples how the Lévy-Min-Sampler generalizes previous sketches and leads to new sketches. First, recall that by Lévy-Khintchine (Theorem˜7), a function $G:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is the Laplace exponent of some subordinator if and only if it is generated by a triplet $(c,\nu,\gamma_{0})$ where $c,\gamma_{0}\geq 0$, and $\int_{0}^{\infty}\min(s,1)\,\nu(ds)<\infty$.
$G(x)$ $=c\mathbbm{1}\!\left\{{x>0}\right\}+\gamma_{0}x+\int_{0}^{\infty}(1-e^{-xs})\,\nu(ds).$
Example 1 ($F_{0}$-sampler $\mapsto$ the Min sketch [13]).
For $F_{0}$-sampling, we have $G(x)=\mathbbm{1}\!\left\{{x>0}\right\}$. which is generated by $(c,\nu,\gamma_{0})=(1,0,0)$. This corresponds to a “pure-killed” process $X$, where, given a kill time $Y\sim\mathrm{Exp}(1)$, $X_{t}=0$ if $t<Y$ and $X_{t}=\infty$ if $t\geq Y$. The induced level function (Definition˜6) is,
$\ell(x,y)$ $=\inf\{z:\mathbb{P}(X_{z}\geq x)\geq y\}$
Since $x>0$, $\mathbb{P}(X_{z}\geq x)=\mathbb{P}(z\geq Y)=1-e^{-z}$. Regardless of $x$, this is equal to
$=\inf\{z:1-e^{-z}\geq y\}$
$=-\log(1-y).$
Thus the resulting sketch stores $\log(1-\eta)$, where $\eta$ is the minimum hash value, thereby essentially reproducing Cohen’s [13] Min sketch.
Example 2 ($F_{1}$-sampler $\mapsto$ min-based reservoir sampling [52]).
For $F_{1}$-sampling, we have $G(x)=x$, which is generated by $(c,\nu,\gamma_{0})=(0,0,1)$. This corresponds to a deterministic drift process $X$, where $X_{t}=t$ for $t\geq 0$. The induced level function is,
$\ell(x,y)$ $=\inf\{z:\mathbb{P}(X_{z}\geq x)\geq y\}$
and since $X_{z}=z$, $\mathbb{P}(X_{z}\geq x)=\mathbbm{1}\!\left\{{z\geq x}\right\}$,
$=\inf\{z:\mathbbm{1}\!\left\{{z\geq x}\right\}\geq y\}$
$=x.$ (note that $y\in(0,1)$)
Thus the resulting sketch stores the minimum random value sampled freshly at each insertion, reproducing the reservoir sampler [52] with the choice of replacement simulated by taking the min.
Next, we demonstrate a new and “non-trivial” application: the construction of an $F_{1/2}$-sampler.
Sketch : $(v_{*},h_{*})$, initialized as $(\perp,\infty)$
Hash function : $H:[n]\to\mathrm{Uniform}(0,1)$
Result: Sample an element $u$ with prob. $\sqrt{\mathbf{x}(u)}/\sum_{v\in[n]}\sqrt{\mathbf{x}(v)}$
// upon update $\mathbf{x}(v)\leftarrow\mathbf{x}(v)+\Delta$
Update $(v,\Delta)$
 $h\leftarrow\sqrt{2\mathrm{Exp}(\Delta)}\cdot\mathrm{erf}^{-1}(H(v))$
 // $\mathrm{Exp}(\Delta)$ is freshly sampled
 if $h<h_{*}$ then
    $(v_{*},h_{*})\leftarrow(v,h)$
   end if
// upon sample
Sample ()(\,)
 return $v_{*}$
Algorithm 3 $F_{1/2}$-sampler
Example 3 ($F_{1/2}$-sampler).
For $F_{1/2}$, we have $G(x)=\sqrt{x}$, which corresponds to the $1/2$-stable process. The induced level function is
$\ell(x,y)$ $=\inf\{z:\mathbb{P}(X_{z}\geq x)\geq y\}$
and since $X$ is $1/2$-stable, we have $X_{z}\sim z^{2}X_{1}$
$=\inf\{z:\mathbb{P}(z^{2}X_{1}\geq x)\geq y\}.$
It is known that the standard 1/2-stable $X_{1}$ distributes identically with $1/Z^{2}$ where $Z$ is a standard Gaussian [50, page 29]. Thus, we have $\mathbb{P}(X_{1}\geq r)=\mathbb{P}(|Z|\leq\sqrt{1/r})=\mathrm{erf}\left(\sqrt{\frac{1}{2r}}\right)$, where $\mathrm{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-x^{2}}\,dx$ is the Gauss error function. As $\mathbb{P}(z^{2}X_{1}\geq x)=\mathbb{P}(X_{1}\geq x/z^{2})=\mathrm{erf}(z/\sqrt{2x})$, this is equal to
$=\sqrt{2x}\cdot\mathrm{erf}^{-1}(y).$
Thus to sample with weight function $G(x)=\sqrt{x}$, upon an $\mathsf{Update}(v,\Delta)$, one just needs to compute $h=\sqrt{2Y/\Delta}\cdot\mathrm{erf}^{-1}(H(v))$ where $Y\sim\mathrm{Exp}(1)$ is freshly sampled and $H(v)\sim\mathrm{Uniform}(0,1)$ is the hash value of $v$, and replace $(v_{*},h_{*})$ with $(v,h)$ if $h<h_{*}$. [^10] The inverse error function $\mathrm{erf}^{-1}$ is available, e.g., as $\mathtt{scipy.special.erfinv}$ in Python. See Algorithm˜3.
The $F_{1/2}$-moment is a special case where the $1/2$-stable distribution has a clean expression. Penson and Górska [46] give explicit formulas for one-sided $k/l$-stable distributions for integers $k<l$. For a generic $\alpha\in(0,1)$ whose level function is difficult to compute, it is more convenient to use the Stable-Min-Sampler (Theorem˜12).
Example 4 (Common subordinators).
We now consider samplers induced by common subordinators.
  • The $F_{\alpha}$-sampler with $G(z)=z^{\alpha}$ for $\alpha\in(0,1)$ corresponds to the non-negative $\alpha$-stable process, where $X_{1}\sim\alpha\text{-stable}$.
  • The sampler with $G(z)=1-e^{-\lambda x}$ for $\lambda>0$ corresponds to the Poisson processes, where $X_{1}\sim\mathrm{Poisson}(\lambda)$.
  • The sampler with $G(z)=\alpha\log(1+x/\beta)$ for $\alpha,\beta>0$ corresponds to Gamma processes, where $X_{1}\sim\mathrm{Gamma}(\alpha,\beta)$.
See Fig.˜2 and Fig.˜3 for contour plots of the level functions used in the examples. For a generic $G$, in practice one may pre-compute the level function of $G$ on a geometrically spaced lattice and cache it as a read-only table. Such a table can be shared and read simultaneously by an unbounded number of $G$-samplers for different applications and therefore the amortized space overhead is typically small.
(a) Level function for $F_{0}$-sampler ($G$-sampler with $G(z)=\mathbbm{1}\!\left\{{z>0}\right\}$)
(b) Level function for $F_{1}$-sampler ($G$-sampler with $G(z)=z$)
Figure 2: Level functions for $F_{0}$-sampler and $F_{1}$-sampler
(a) Level function for $(1-e^{-z})$-sampler, induced by a Poisson counting process
(b) Level function for $z^{1/2}$-sampler (a.k.a $F_{1/2}$-sampler), induced by a $1/2$-stable process
(c) Level function for $\log(1+z)$-sampler, induced by a Gamma process
Figure 3: Level functions for samplers induced by well-known Lévy processes
4.4 Universal Sampler
Since $\ell_{G}(a,b)$ is increasing in both arguments, among all updates $\{(v_{i},\Delta_{i})\}$ to the Lévy-Min-Sampler, the stored sample must correspond to a point on the (minimum) Pareto frontier of $\{(Y_{i}/\Delta_{i},H(v_{i}))\}$. Thus, it is possible to produce a $G$-sample for any $G\in\mathcal{G}$ simply by storing the Pareto frontier. (This observation was also used by Cohen [15] in her approximate samplers.) The size of the Pareto frontier is a random variable that is less than $\ln n+1$ in expectation and $O(\log n)$ with high probability.
Theorem 14.
Suppose ParetoSampler processes a stream of $\operatorname{poly}(n)$ updates to $\mathbf{x}$. The maximum space used is $O(\log n)$ words with probability $1-1/\operatorname{poly}(n)$. At any time, given a $G\in\mathcal{G}$, it can produce a $v_{*}\in[n]$ such that $\mathbb{P}(v_{*}=v)=G(\mathbf{x}(v))/G(\mathbf{x})$.
By instantiating $k$ independent copies of the Lévy-Min-Sampler we can $G$-sample $k$ indices with replacement, but in some applications we would like to sample $k$ indices without replacement. In Appendix A, we show that both Lévy-Min-Sampler and ParetoSampler can be modified to sample without replacement as well.
Sketch : $S\subset\mathbb{R}_{+}\times\mathbb{R}_{+}\times[n]$, initialized as $\emptyset$
Hash function : $H:[n]\to\mathrm{Uniform}(0,1)$
Result: Sample an element $u$ with prob. $G(\mathbf{x}(u))/\sum_{v\in[n]}G(\mathbf{x}(v))$ upon a specified $G$
// upon update $\mathbf{x}(v)\leftarrow\mathbf{x}(v)+\Delta$
Update $(v,\Delta)$
 $S\leftarrow\textsf{Pareto}(S\cup\{(\mathrm{Exp}(\Delta),H(v),v)\})$
 // $\mathrm{Exp}(\Delta)$ is freshly sampled
// upon $G$-sample
Sample $(G)$
 $(a_{*},b_{*},v_{*})\leftarrow\operatorname{argmin}_{(a,b,v)\in S}\{\ell_{G}(a,b)\}$
 return $v_{*}$
Algorithm 4 ParetoSampler. The function $\textsf{Pareto}(L)$ returns the (minimum) Pareto frontier of the tuples $L$ w.r.t. their first two coordinates.
5 The LévyPCSA and LévyHyperLogLog Sketches
The $G$-transformation technique can be used to reduce $G$-moment estimation in the $\mathbb{R}_{+}$-turnstile to the intensively studied cardinality estimation problem.
5.1 Lévy-Activation and $G$-Cells
Ting [51] and Pettie and Wang [47] analyze cardinality sketches in a uniform fashion by viewing them as a set $\mathcal{C}$ of cells which begin inactive (0) and can make a one-time transition to active (1). The probability that an update activates a cell $c\in\mathcal{C}$ is proportional to its size $s(c)$. Without loss of generality, one may assume the sizes are normalized to that $\sum_{c\in\mathcal{C}}s(c)=1$. The state of the sketch is the set of active cells. See [47, §1.2] for more details.
Definition 7 (activation of cells).
Let $H:[n]\times\mathcal{C}\to[0,1]$ be a uniformly random hash function. Let $s(c)$ be the size of cell $c\in\mathcal{C}$. Initially $c\leftarrow 0$ for all $c\in\mathcal{C}$. Upon an update with $\Delta>0$,
$\mathsf{Update}(v,\Delta)$ :
For each $\in\mathcal{C}, cc$$\leftarrow c\vee\mathbbm{1}\!\left\{{H(v,c)<s(c)}\right\}$.
Assume the total cardinality is $\lambda\in\mathbb{N}$. A cell with size $s$ will be activated with probability $1-(1-s)^{\lambda}\to 1-e^{-s\lambda}$ as $\lambda\to\infty$ and $s\lambda=O(1)$. All upper bounds for cardinality estimation in the random oracle model can be applied to the $G$-moment estimation if cells with activation probability $1-e^{-sG(\mathbf{x})}$ can be simulated. We call the simulated cells $G$-cells.
Definition 8 (Lévy-activation of $G$-cells).
Let $X$ be a subordinator and $G$ be the corresponding Laplace exponent. Upon an update with $\Delta>0$,
$\mathsf{Update}(v,\Delta)$:
For each $\in\mathcal{C}, cc$$\leftarrow c\vee\mathbbm{1}\!\left\{{X^{(v,c)}_{s(c)}>Y/\Delta}\right\}$, where $X^{(v,c)}\sim X$ are i.i.d. copies of $X$ and $Y\sim\mathrm{Exp}(1)$ is a freshly sampled standard exponential random variable.
We now prove that the $G$-cells simulate the ordinary cells in Definition˜7 asymptotically with the cardinality $\lambda$ replaced by the $G$-moment.
Lemma 9.
Given any input vector $\mathbf{x}\in\mathbb{R}_{+}^{n}$, the activation of $G$-cells are independent and for each $G$-cell $c$,
$\mathbb{P}(c=1)$ $=1-e^{-s(c)G(\mathbf{x})}.$
Proof.
The independence is given by construction.
$\mathbb{P}(c=0)$ $=\mathbb{P}\left(\forall v\in[n].\;X_{s(c)}^{(v,c)}\leq\mathrm{Exp}(\mathbf{x}(v))\right)$
$=\prod_{v\in[n]}\mathbb{P}\left(X_{s(c)}^{(v,c)}\leq\mathrm{Exp}(\mathbf{x}(v))\right)$ (by independence)
$=\prod_{v\in[n]}\mathbb{E}e^{-X_{s(c)}^{(v,c)}\mathbf{x}(v)}$
$=\prod_{v\in[n]}e^{-s(c)G(\mathbf{x}(v))}=e^{-s(c)G(\mathbf{x})}.$ (by Lévy-Khintchine)
Hence $\mathbb{P}(c=1)=1-e^{-s(c)G(\mathbf{x})}$. ∎
The Lévy activation lemma enables essentially all cardinality estimation techniques to be applied to $G$-moment estimation, by replacing cells with $G$-cells. We first consider the LévyPCSA sketch, i.e., the sketch that replaces every cell in PCSA with a $G$-cell with the same size.111111(In brief, the basic PCSA sketch has cells with sizes $1/2,1/4,1/8,\ldots$. A PCSA sketch composed of $m$ subsketches has $m$ cells of size $m^{-1}\cdot 2^{-i}$ for all $i\geq 1$. See [24, 47, 51] for more details.)
Theorem 15 (PCSA emulation).
Given any input vectors $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}_{+}^{n}$ such that $G(\mathbf{x})=\|\mathbf{x}^{\prime}\|_{0}$, LévyPCSA with input $\mathbf{x}$ and Poissonized PCSA with input $\mathbf{x}^{\prime}$ distribute identically.
The most efficient mergeable cardinality sketch is Fishmonger [47], which is an entropy-compressed version of PCSA with an asymptotically optimum estimator. It is optimal among all “linearizable” mergeable sketches [47]. (Lang’s CPC sketch [38] included in Apache DataSketches [3] is similar, but uses an off-the-shelf compressor that does not meet the entropy bound and uses an estimator [38] that is worse than [55, 47].) Substituting cells with $G$-cells in Fishmonger yields LévyFishmonger.
Corollary 1.
The LévyFishmonger sketch [47] provides an asymptotically unbiased estimate of the $G$-moment with relative variance $1/m$. It uses $(1+o(1))(H_{0}/I_{0})m\approx 1.98m$ bits when $m=\Omega(\log^{2}\log n)$. Here $H_{0}=\frac{1}{\log 2}+\sum_{k=1}^{\infty}\frac{1}{k}\log_{2}\left(1+1/k\right)$, $I_{0}=\pi^{2}/6$, and $H_{0}/I_{0}\approx 1.98$.
Although (compressed) PCSA has the best size-accuracy tradeoff [47, 38, 3], HyperLogLog is already widely deployed in industry. For these reasons the LévyHyperLogLog sketch will be more attractive in some real world applications. Let $H:[n]\to\mathbb{N}$ be a hash function where $\mathbb{P}(H(v)\geq k)=2^{-k}$ for $k\in\mathbb{N}$. One HyperLogLog subsketch [23] stores a single number $M$ such that upon insertion $v$, $M\leftarrow\max(M,H(v))$. With $\mathrm{Poisson}(\lambda)$ distinct insertions, $\mathbb{P}(M\geq k)=1-e^{-\lambda 2^{-k}}$. We now show how to emulate $M$ with $\lambda$ replaced by $G(\mathbf{x})$.
Lemma 10 (LévyHyperLogLog).
Let $G$ be the Laplace exponent and $X$ be the corresponding subordinator. Let $M$ be an $\mathbb{N}$-register initialized as 0. A vector update $\mathbf{x}(v)\leftarrow\mathbf{x}(v)+\Delta$ is effected by:
$\mathsf{Update}(v,\Delta)$ :
$M\leftarrow\max\left\{M,\max\left\{k\in\mathbb{N}:X^{(u)}_{2^{-k}}>Y/\Delta\right\}\right\}$, where $X^{(v)}\sim X$ is hashed with key $v$ and $Y\sim\mathrm{Exp}(1)$ is a freshly sampled standard exponential random variable.
Then given any input vector $\mathbf{x}\in\mathbb{R}_{+}^{n}$, $\mathbb{P}(M\geq k)=1-e^{-2^{-k}G(\mathbf{x})}$.
Remark 5.
HyperLogLog is restored by taking $G(x)=\mathbbm{1}\!\left\{{x>0}\right\}$, with $X$ being the “pure killed” process.
Proof.
For any $u\in[n],\mathbf{x}(u)>0$, and $j\in\mathbb{N}$,
$\mathbb{P}\left(\max\left\{k\in\mathbb{N}:X^{(u)}_{2^{-k}}>\mathrm{Exp}(\mathbf{x}(u))\right\}<j\right)$
$=\mathbb{P}\left(X^{(u)}_{2^{-j}}\leq\mathrm{Exp}(\mathbf{x}(u))\right)$
$=\mathbb{E}e^{-\mathbf{x}(u)X^{(u)}_{2^{-j}}}$
$=e^{-2^{-j}G(\mathbf{x}(u))}.$ (by Lévy-Khintchine)
Thus
$\mathbb{P}(M<j)$ $=\prod_{u\in[n]}\mathbb{P}\left(\max\left\{k\in\mathbb{N}:X^{(u)}_{2^{-k}}>\mathrm{Exp}(\mathbf{x}(u))\right\}<j\right)\>=\>e^{-2^{-j}\sum_{u\in[n]}G(\mathbf{x}(u))}=e^{-2^{-j}G(\mathbf{x})}.$
Theorem 16 (HyperLogLog emulation).
Given any input vector $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}_{+}^{n}$ such that $G(\mathbf{x})=\|\mathbf{x}^{\prime}\|_{0}$, LévyHyperLogLog with input $\mathbf{x}$ and Poissonized HyperLogLog with input $\mathbf{x}^{\prime}$ distribute identically.
This allows us to easily adapt any existing implementation of HyperLogLog to estimate $G$-moments, using any estimator for HyperLogLog [23, 55, 22].
5.2 Stable-HyperLogLog
For $\alpha$-stable processes, it is sufficient to hash every element to a single one-sided $\alpha$-stable random variable, instead of an $\alpha$-stable sample path. Lemma˜11 establishes the correctness of Algorithm˜5, whose Query algorithm includes both Flajolet et al.’s [23] harmonic mean estimator and Wang and Pettie’s [55] improved $\tau$-GRA-based estimator.
Sketch : $M[1],\ldots,M[m]$, initialized as -$\infty$
Hash function :  $H:[n]\times[m]\to\textrm{one sided $\alpha$-stable}$, $\alpha\in(0,1)$
Result: Estimate $\sum_{v\in[n]}\mathbf{x}(v)^{\alpha}$
// upon update $\mathbf{x}(v)\leftarrow\mathbf{x}(v)+\Delta$
Update $(v,\Delta)$
 for $j\in[m]$ do
    $w\leftarrow\lfloor\alpha(\log_{2}H(v,j)-\log_{2}\mathrm{Exp}(\Delta))\rfloor$
    // $\mathrm{Exp}(\Delta)$ is freshly sampled
    $M[j]\leftarrow\max(M[j],w)$
   end for
// upon query
Query ()(\,)
 return $C_{m}(\frac{1}{m}\sum_{j=1}^{m}2^{-M[j]})^{-1}$
 // HyperLogLog’s estimator; $C_{m}\to 0.7213$
   or return $(\frac{\log 2}{\Gamma(\tau_{*})(1-2^{-\tau_{*}})}\frac{1}{m}\sum_{j=1}^{m}2^{-\tau_{*}M[j]})^{-\tau_{*}^{-1}}$
 // $\tau_{*}$-GRA estimator; $\tau_{*}\to 0.89$
Algorithm 5 $F_{\alpha}$-moment estimation (Stable-HyperLogLog)
Lemma 11 (Stable-HyperLogLog).
Fix $\alpha\in(0,1)$. Let $M$ be an $\mathbb{N}$-register initialized as 0. A vector update $\mathbf{x}(v)\leftarrow\mathbf{x}(v)+\Delta$, $\Delta>0$, is effected by:
$M\leftarrow\max\left\{M,\left\lfloor\alpha(\log_{2}Z^{(v)}-\log_{2}Y+\log_{2}\Delta)\right\rfloor\right\},$
where the $Z^{(v)}$ are i.i.d. standard one-sided $\alpha$-stable random variables and $Y\sim\mathrm{Exp}(1)$ is a freshly sampled standard exponential random variable. For any input vector $\mathbf{x}\in\mathbb{R}_{+}^{n}$, $\mathbb{P}(M\geq k)=1-e^{-2^{-k}\sum_{v\in[n]}\mathbf{x}(v)^{\alpha}}$.
Remark 6.
For comparison, note that at $\mathsf{Update}(v,\Delta)$, for any $\Delta>0$, HyperLogLog updates as
$M\leftarrow\max(M,\lfloor-\log_{2}W^{(v)}\rfloor),$
where $W^{(v)}\sim\mathrm{Exp}(1)$. HyperLogLog can be considered as the limiting case of Stable-HyperLogLog, since
$\alpha(\log_{2}Z^{(v)}-\log_{2}Y+\log_{2}\Delta)\to-\log_{2}W^{(v)}$
in distribution as $\alpha\to 0$. It is known that $(\textrm{$\alpha$-stable})^{\alpha}\to 1/\mathrm{Exp}(1)$ in distribution as $\alpha\to 0$ [20].
Proof.
For any $u\in[n],\mathbf{x}(u)>0$, and $j\in\mathbb{N}$,
$\mathbb{P}\left(\lfloor\alpha(\log_{2}Z^{(u)}-\log_{2}\mathrm{Exp}(\mathbf{x}(u)))\rfloor<j\right)$
$=\mathbb{P}\left(\max\left\{k\in\mathbb{N}:Z^{(u)}>2^{k/\alpha}\mathrm{Exp}(\mathbf{x}(u))\right\}<j\right)$
$=\mathbb{P}\left(Z^{(u)}\leq 2^{j/\alpha}\mathrm{Exp}(\mathbf{x}(u))\right)$
$=\mathbb{E}e^{-\mathbf{x}(u)2^{-j/\alpha}Z^{(u)}}.$
By Lévy-Khintchine, $\mathbb{E}e^{-zZ^{(u)}}=e^{-z^{\alpha}}$ for $z\geq 0$ since $Z^{(u)}$ is $\alpha$-stable, so this is equal to
$=e^{-2^{-j}\mathbf{x}(u)^{\alpha}}.$
It follows that $\mathbb{P}(M<k)=\prod_{u\in[n]}\mathbb{P}\left(\lfloor\alpha(\log_{2}Z^{(u)}-\log_{2}\mathrm{Exp}(\mathbf{x}(u)))\rfloor<k\right)=2^{-2^{-k}\sum_{u\in[n]}\mathbf{x}(u)^{\alpha}}$. ∎
6 Previous Sketches and Lévy Processes
In this section we will discuss in greater detail how previous sketches can be viewed as being constructed from Lévy processes. The guiding question is: what will happen if the input vector is replicated a large number of times on disjoint supports?
6.1 Single-Level Aggregation and the Central Limit Theorem
We start from the AMS sketch by Alon, Matias, and Szegedy [1], where a cell $Q$ stores
$Q$ $=\sum_{v\in[n]}\mathbf{x}(v)\xi_{v}.$
Here $\mathbf{x}\in\mathbb{R}^{n}$ is updated in an $\mathbb{R}$-turnstile and the $\xi_{v}\in\{-1,1\}$ are i.i.d. Rademacher random variables. [^12] Again, in Alon, Matias, and Szegedy’s analysis [1], 4-wise independent hashing suffices to guarantee a good enough estimate but we need to assume they are i.i.d. here to talk about the exact distribution of the final state. Now suppose the input vector is repeated $w$ times on disjoint supports, i.e., $\mathbf{x}_{w}\in\mathbb{R}^{nw}$ with $\mathbf{x}_{w}(v)=\mathbf{x}(v\bmod n)$. Then the final state of the sketch for $\mathbf{x}_{w}$ would be
$M_{w}=Q_{1}+Q_{2}+\cdots+Q_{w},$
where the $(Q_{j})$ are i.i.d. copies of $Q$. Note that $\mathbb{E}Q=0$ and $\mathbb{V}Q=|\mathbf{x}|^{2}=\sum_{v\in[n]}\mathbf{x}(v)^{2}$. Thus as $w\to\infty$ the normalized final state $M_{w}/\sqrt{w}$ converges to a centered Gaussian random variable with variance $\sum_{v\in[n]}\mathbf{x}(v)^{2}$ in distribution, by the central limit theorem. Thus, in the regime as $w\to\infty$, the AMS sketch eventually becomes the same as a cell in the Lévy-Tower $\sum_{v\in[n]}\left\langle X^{(v)}_{1},\mathbf{x}(v)\right\rangle$, where the $(X^{(v)})$ are i.i.d. one-dimensional Wiener processes/Brownian motion and $X_{1}^{(v)}$ is the value of the process at time $t=1$. Of course, since a Wiener process/Brownian motion is self-similar, one does not need to sample the Lévy processes at exponentially spaced intervals; samples from $m$ independent processes at time $t=1$ suffice to approximate $\sum_{v\in[n]}\mathbf{x}(v)^{2}$ with $O(1/m)$ relative variance. The characteristic exponent of the Wiener process/Brownian motion is the target function for the $F_{2}$-moment: $f(x)=|x|^{2}$.
Another illuminating example is Ganguly’s [25] $F_{k}$-moment estimator for $k\geq 3$. Ganguly randomly projects the elements and stores $S=\sum_{v\in[n]}\mathbf{x}(v)Z^{(v)}$ where the $Z^{(v)}$ are i.i.d. uniformly random roots of $x^{k}=1$ on the complex unit circle. The statistic $S^{k}$ seems to be a good estimator for the $F_{k}$-moment since $\mathbb{E}S^{k}=\sum_{v\in[n]}\mathbf{x}(v)^{k}$. However, since this random projection has finite variance, the normalized sum $S/\sqrt{\sum_{v\in[n]}|\mathbf{x}(v)|^{2}}$ will converge to a complex Gaussian as the input vector is replicated on disjoint supports. In the limit, the only information remaining in the sketch pertains to the $F_{2}$-moment. Indeed, one in fact needs the number of i.i.d. registers to grow polynomially in the support-size for it to be able to estimate $F_{k}$ [25].
By the generalized central limit theorem, if for some sequences $(a_{w})_{w\in\mathbb{N}}$ and $(b_{w})_{w\in\mathbb{N}}$, $(M_{w}-b_{w})/a_{w}$ converges to some non-degenerate random variable $Y$ as $w\to\infty$, then $Y$ has to be $\alpha$-stable for $\alpha\in(0,2]$. If in addition, $\mathbb{V}Q<\infty$, then $Y$ is Gaussian. Non-Gaussian stable distributions will be discussed shortly, in Section˜6.3.
6.2 Multi-Level Subsampling and the Poisson Limit Theorem
Another important sketching technique is subsampling. The idea is to devise a sketch that works for $\Theta(m)$ elements and then solve the generic case by subsampling the stream at rates $2^{-k}$ for $k\in\mathbb{N}$, one of which reduces it down to $\Theta(m)$ elements. Without loss of generality, suppose now we have $m$ non-zero elements with distinct values $\mathbf{x}(1),\ldots,\mathbf{x}(m)$. Once again, let $\mathbf{x}_{w}\in\mathbb{R}^{nw}$ be $\mathbf{x}$ repeated $w$ times on disjoint supports. To obtain a $\Theta(m)$-size set of subsamples, one needs to subsample $\mathbf{x}_{w}$ with rate $1/w$. Let $Y_{w,j}$ be the indicator that the $j$th copy of $\mathbf{x}(1)$ is sampled where $\mathbb{E}Y_{w,j}=1/w$. The number of elements with value $\mathbf{x}(1)$ in the subsampled set is
$\sum_{j=1}^{w}Y_{w,j}\to\mathrm{Poisson}(1)\text{ in distribution as $w\to\infty$}.$
Here we have invoked the Poisson limit theorem (see [21, Theorem 3.6.1]) which can be applied since $\mathbb{E}\sum_{j=1}^{w}Y_{w,j}=1$ and $\max_{j=1}^{w}\mathbb{P}(Y_{w,j}\neq 0)=\frac{1}{w}\to 0$ as $w\to\infty$. Similarly, the number of elements with value $\mathbf{x}(j)$ is also $\mathrm{Poisson}(1)$ and the occurrences of different values are independent. It is well known that such a limiting distribution can be simulated algorithmically by duplicating every element a $\mathrm{Poisson}(1)$ number of times (see [24, 23, 47, 55]). On the other hand, such limit distributions can be equivalently simulated by hashing each element to a Poisson process and then sampling at different times. Thus, all sketches based on subsampling can be simulated by the Lévy-Tower with the corresponding (compound) Poisson processes.
6.3 Stable Random Variables and Stable Processes
Whenever $Q$ has a finite variance, $(Q_{1}+Q_{2}+\cdots+Q_{w})/\sqrt{w}$ goes to Gaussian as $w\to\infty$. Nevertheless, with infinite variance, $Q$ can lie in the domain of attraction of an $\alpha$-stable distribution for any $\alpha\in(0,2)$. Indeed, the use of stable random variables is another sketching technique that directly corresponds to Lévy processes, namely $\alpha$-stable processes. Similar to the Gaussian case (Section˜6.1), $\alpha$-stable processes are self-similar so there is no need to store the whole tower with exponentially spaced sample times $2^{-k}$. Rather, it suffices to sample $m$ independent processes at time $t=1$.
Indyk [29] uses one-dimensional $\alpha$-stable random variables for $\alpha\in(0,2]$ to estimate the $F_{\alpha}$-moment, and indeed, the characteristic exponent of the $\alpha$-stable process is $f(x)=|x|^{\alpha}$. Ganguly, Bansal, and Dube [28] consider the higher dimensional case, where each element has $d$ attributes. We now show how to reconstruct [28]’s sketch from the perspective of Lévy processes. The target function is
$f:\mathbb{R}^{d}\to\mathbb{R},\quad f(x)$ $=\left(\sum_{j=1}^{d}|x_{j}|^{p}\right)^{q},$
where $p\in(0,2],q\in(0,1]$. Such $f$-moments are called $F_{p,q}$ hybrid moments in [28]. By the subordination theorem (Theorem˜8; see [50, page 197]), this function is the characteristic exponent of a vector of $d$ independent $p$-stable processes that are subordinated by a common $q$-stable subordinator. That is, the Lévy process is precisely $Y_{t}\in\mathbb{R}^{d}$,
$Y_{t}=(X_{Z_{t}}^{[1]},\ldots,X_{Z_{t}}^{[d]}),$
where $Z$ is a $q$-stable subordinator and the $(X^{[j]})$ are i.i.d. $p$-stable processes. Note that since $X^{[j]}$ is $p$-stable, we have $X_{Z_{t}}^{[j]}\sim Z_{t}^{1/p}X_{1}^{[j]}$. Thus, sampling $Y$ at time 1 yields
$Y_{1}\sim Z_{1}^{1/p}\left(X_{1}^{[1]},\ldots,X_{1}^{[d]}\right).$
Take the inner product and we have, for $x\in\mathbb{R}^{d}$,
$\left\langle x,Y_{1}\right\rangle=Z_{1}^{1/p}\sum_{j=1}^{d}x_{j}X_{1}^{[j]},$
which is exactly the random projection defined by the update algorithm of Ganguly, Bansal, and Dube [28].
Remark 7.
One benefit from understanding the connection between moment estimation and Lévy processes is debugging incorrect claims. The conference version of Ganguly, Bansal, and Dube [27] stated that the $F_{p,q}$ hybrid moment could be estimated in $\operatorname{polylog}(n)$ space for any $p,q\in(0,2]$. This claim was shown to be incorrect by Jayram and Woodruff [34]. Using the connection between Lévy processes and sketches, this claim should seem fishy, as the $F_{p,q}$-hybrid moment corresponds to a Lévy process consisting of a vector of $p$-stable $X^{[1]},\ldots,X^{[d]}$ subordinated by a common $q$-stable subordinator. But $q$-stable subordinators do not exist for $q\in(1,2]$! This error was corrected in the journal version of Ganguly, Bansal, and Dube [28], where the space is $\operatorname{polylog}(n)$ for $p\in(0,2],q\in(0,1]$ and polynomial for $p\in(0,2],q\in(1,2]$. (See Jayram and Woodruff [34] for $F_{p,q}$-moment estimators for general $p,q$.)
6.4 HyperLogLog, PCSA, and Pure Killed Processes
Cardinality sketches like HyperLogLog [23] and PCSA [24] that only allow increments also correspond to Lévy processes in a surprisingly natural way. We consider the number system $\mathbb{R}\cup\{\infty\}$, where for any $\in\mathbb{R}, x+=+x$$\infty=\infty+x$$\infty$ and $\infty+\infty=\infty$. Such definitions extend to multiplication with natural numbers where $0\cdot\infty=0$ and $k\cdot\infty=\infty$ for any $k\in\mathbb{Z}_{+}$.
Definition 9 (Pure killed processes).
A pure killed process $X=(X_{t})_{t\geq 0}$ with kill rate $c>0$ is a Lévy process over $\mathbb{R}\cup\{\infty\}$ which can be simulated as follows.
  • Sample a kill time $Y\sim\mathrm{Exp}(c)$.
  • $X_{t}=0$ if $t<Y$ and $X_{t}=\infty$ otherwise.
In particular, we have $\mathbb{P}(X_{t}=\infty)=\mathbb{P}(Y\leq t)=1-e^{-ct}$.
Assume the insertion stream is $v_{1},\ldots,v_{T}$, where each $v_{j}$ is an element in the universe $[n]$. Let $X=(X_{t})_{t\geq 0}$ be a pure killed process with unit kill rate $(c=1)$. Similar to the turnstile case in Section˜1.3, one stores the sum at time $t$,
$C_{t}$ $=\sum_{j=1}^{T}X_{t}^{(v_{j})}=\sum_{v\in[n]}\mathbf{x}(v)\cdot X_{t}^{(v)}\in\{0,\infty\},$
where $X^{(v)}$s are i.i.d. copies of $X$. It is straightforward to see that $S_{k}$ perfectly simulates a Poissonized PCSA cell [24, 47], with $(\{0,\infty\},+)$ mapped to $(\{0,1\},\vee)$. See Table˜3.
Bit operation in cardinality sketches Jumps of a pure killed process
$\lor 0=0 0+0$=00+0=0
$\lor 1=1 0+=0$$\infty$$\infty$
$1\lor 0=1$ $\infty+0=\infty$
$1\lor 1=1$ $\infty+\infty=\infty$
Table 3: HyperLogLog and PCSA can be considered as Lévy-Tower sketches with pure killed processes where the bit-or operation $(\{0,1\},\lor)$ is simulated with the extended real numbers $(\{0,\infty\},+)$. Note that, of course, such a reinterpretation does not upgrade HyperLogLog and PCSA to work over turnstile streams. The resulting Lévy-Tower is still incremental-only since $\infty$-$\infty$ is not defined.
Since $X_{t}^{(v)}\in\{0,\infty\}$, it only matters whether $x(v)$ is zero or non-zero, and we have
$\mathbb{E}e^{iC_{t}}$ $=\prod_{v\in[n]}\mathbb{E}e^{i\mathbf{x}(v)\cdot X_{t}^{(v)}}=\left(\mathbb{E}e^{iX_{t}}\right)^{\|\mathbf{x}\|_{0}}.$ (10)
By Definition˜9, $X_{t}=0$ with probability $e^{-t}$, in which case $e^{iX_{t}}=1$, and $X_{t}=\infty$ with probability $1-e^{-t}$, in which case $e^{iX_{t}}=0$. [^13] Here we take the Cesàro limit $e^{i\infty}=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}e^{it}\,dt=0$. Thus we have $\mathbb{E}e^{iX_{t}}=e^{-t}$. Inserting this back to Eq.˜10, we have
$\mathbb{E}e^{iC_{t}}$ $=e^{-t\|\mathbf{x}\|_{0}}.$
This coincides with the observation in Section˜1.3, where the estimation target $\|\mathbf{x}\|_{0}$ lies right in the exponent of $\mathbb{E}e^{iC_{t}}$.
[…]
6.5 Uniform Random Projection and the Integral Lévy-Tower
Indyk’s $F_{p}$-stable sketches [29] are able to estimate the $F_{p}$-moment for $p\in(0,2]$ but do not handle the $F_{0}$-moment. Cormode, Datar, Indyk, and Muthukrishnan [18] approximate the $F_{0}$-moment by the $F_{\alpha}$-moment for very small $\alpha>0$. This scheme needs a bound on $\|\mathbf{x}\|_{\infty}$, as a single unbounded update can raise the estimate to infinity. Kane, Nelson, and Woodruff [37] take another approach: subsample the stream to a suitable level, randomly project each element within the field $\mathbb{Z}_{p}$ for a random prime $p>\epsilon^{-1}\log(\|\mathbf{x}\|_{\infty})$, and then use an (insertion-only) cardinality estimator. The random projection approach [37] also requires a bound on $\|\mathbf{x}\|_{\infty}$. As an $F_{0}$-moment sketch it has a certain failure probability (e.g., if $p|\mathbf{x}(v)$ for many $v$) but it can be viewed as estimating a related quantity: the $f_{F_{0},p}$-moment for $f_{F_{0},p}(x)=\mathbbm{1}\!\left\{{p\nmid x}\right\}$.
We show how the Lévy-Tower reconstructs Kane, Nelson, and Woodruff’s [37] trick mechanically, by pure computation. The target function $f(x)=\mathbbm{1}\!\left\{{p\nmid x}\right\}$ for $x\in\mathbb{Z}$ can be written as the following Lévy-Khintchine-representable function; see Remark˜1.
$f(x)$ $=\frac{1}{p}\sum_{j=0}^{p-1}(1-\cos(2\pi jx/p)).$
When $p\mid x$ each of the summands is 0 whereas when $p\nmid x$ the cosine terms sum to 0. The Lévy process with characteristic exponent $f$ is the compound Poisson process with jumps uniformly chosen from $2\pi(p-1)/p,\ldots,0,\ldots,2\pi(p-1)/p$. See Fig.˜4. The corresponding Lévy-Tower reconstructs the Kane-Nelson-Woodruff sketch [37]: subsample through Poisson processes and uniformly project over $\mathbb{Z}_{p}$ with the uniform random jumps. It is straightforward to see that the resulting Lévy-Tower is integral in the sense that $S_{k}\in\{2j\pi/p:j=0,\ldots,p-1\}$ so $S_{k}$ can be identified by a $\mathbb{Z}_{p}$-value.
Figure 4: Left: $f_{F_{0},32}(x)=\sum_{j=1}^{31}\frac{1}{32}(1-\cos(2\pi jx/32)).$ The black dots mark the values at integer $x$s. Right: The jump distribution of the compound Poisson process $X$ with characteristic exponent $f_{F_{0},32}$. The subsampling and uniform random projection tricks used in [37] are recovered from computing the corresponding Lévy process.
7 Tractability and Lévy Processes
The most basic complexity-theoretic question in any model of computation is to separate the tractable from the intractable, e.g., to characterize the class $\mathbf{P}$ in Turing-machine complexity. Braverman and Ostrovsky [9] defined an analogue of $\mathbf{P}$ for $f$-moment estimation.
Definition 10 (Tractability).
A function $f:\mathbb{Z}\to\mathbb{R}$ is tractable if it is possible to $(1\pm\epsilon)$-approximate the $f$-moment of a vector $\mathbf{x}\in\{-n,\ldots,0,\ldots,n\}^{n}$ updated in a turnstile stream, with a $\operatorname{poly}(\epsilon^{-1},\log n)$-bit sketch.
Braverman, Chestnut, Woodruff, and Yang [8] relaxed this definition to include $n^{o(1)}\operatorname{poly}(\epsilon^{-1})$-bit sketches, which slightly broadens the class of tractable functions, but does not seem to affect its character very much. Braverman and Ostrovsky [9] characterized the set of tractable $f$ that are monotonically non-decreasing on $0,\infty)$, and Braverman, Chestnut, Woodruff, and Yang [8] gave a nearly complete characterization of all tractable functions—monotone or not—excluding what they termed “nearly periodic functions.” As our results apply to $f$-moments where $f:\mathbb{R}\to\mathbb{C}$ and $f:\mathbb{R}^{d}\to\mathbb{C}$, we expand the definition of tractability to explicitly include the domain.
Definition 11 ($M$-Tractability).
Let $M$ be a commutative monoid. A function $f:M\to\mathbb{C}$ is tractable if for any parameter $\epsilon>0$, it is possible to $(1\pm\epsilon)$-approximate the $f$-moment of a vector $\mathbf{x}\in M^{n}$ updated in an $M$-turnstile stream (Definition˜1), with a $\operatorname{poly}(\epsilon^{-1},\log n)$-bit sketch. (When $M$ is discrete, say $\mathbb{Z}$, we must assume some bound $\|\mathbf{x}\|_{\infty}<n$. When $M$ is continuous let us additionally suppose that there are $\operatorname{poly}(n)$ updates, and each update is presented to $O(\log n)$ bits of precision.) Define $\mathcal{T}[M]$ to be the set of all $M$-tractable functions.
The present work gives us a completely new way to approach the tractability question. We have shown that the Lévy-Tower estimates every function $f$ that is the characteristic exponent of a Lévy process, and such exponents are characterized by the Lévy-Khintchine representation theorem. In this section we investigate the class of tractable functions captured by the Braverman et al. [8] sketch and the Lévy-Khintchine-representable functions captured by the Lévy-Tower sketch. The Lévy-Khintchine-class includes many multidimensional functions $f:\mathbb{R}^{d}\to\mathbb{C}$ outside the scope of Braverman et al. [8], so we shall mainly investigate each classes’s coverage of $\mathcal{T}[\mathbb{Z}]$ and implications for $\mathcal{T}[\mathbb{R}]$. In the remainder of this section we present the following observations and results.
I.
We show that some easy tests for intractability have a basis in non-Lévy-Khintchine-representable functions.
II.
We prove that a function can be $\mathbb{Z}$-tractable but not $\mathbb{R}$-tractable, that is, $\mathcal{T}[\mathbb{Z}]$ is a strict superset of $\mathcal{T}[\mathbb{R}]$.
III.
We demonstrate that the class of tractable functions implied by the Lévy-Tower/Lévy-Khintchine representation theorem includes nearly periodic functions not captured by the $F_{2}$-heavy hitter framework of [9, 8].
IV.
We observe that there exist simple, tractable functions that do not correspond to the characteristic exponent of any Lévy process. Nonetheless, we can prove that such functions are tractable in the Lévy-Tower/Lévy-Khintchine-framework via what we call the Fourier-Hahn-Lévy method. We conjecture that this method is powerful enough to capture all $M$-tractable moments, for $M\in\{\mathbb{Z},\mathbb{R},\mathbb{Z}^{d},\mathbb{R}^{d}\}$.
7.1 I: Intractability and Non-Lévy-Khintchine-representability
We give a comparison between one-dimensional real characteristic exponents and existing sketching lower bounds based on communication complexity lower bounds. Such exponents can be written as $f(x)=Ax^{2}+\int_{0}^{\infty}(1-\cos(xs))\,\nu(ds)$, since $f$ being real implies that $\nu$ is symmetric; see Remark˜1.
Lemma 12.
Let $f(x)=Ax^{2}+\int_{0}^{\infty}(1-\cos(xs))\,\nu(ds)$. The following statements are true.
  • $f(x)\geq 0$ for any $x\in\mathbb{R}$. [Commentary: functions with both positive and negative values require $\Omega(\operatorname{poly}(n))$-size sketches for constant factor approximations [12].]
  • For any $z\in\mathbb{Z}$ and $x\in\mathbb{R}$, $f(zx)\leq z^{2}f(x)$. [Commentary: functions increasing faster than quadratic require $\Omega(\operatorname{poly}(n))$-size sketches for constant factor approxiamtions [1, 4].]
  • For any $z\in\mathbb{R}_{+}$, $f(z)=0$ if and only if $f(x+z)=f(x),\forall x\in\mathbb{R}$, i.e., $f$ is periodic with period $z$. [Commentary: functions that have zeros other than the origin require $\Omega(\operatorname{poly}(n))$-size sketches, unless the function is periodic with period $\min\{z>0:f(z)=0\}$ [12].]
Proof.
The first statement is obvious. For the second statement, it suffices to prove that with $y=xs$,
$(1-\cos(zy))\leq z^{2}(1-\cos(y)),$
for any $y\in\mathbb{R},z\in\mathbb{Z}$. When $z=0$, the statement trivially holds. Without loss of generality, assume $z\in\mathbb{Z}_{+}$ and let $g(y)=z^{2}(1-\cos(y))-(1-\cos(zy))$. Check that $g^{\prime}(y)=z^{2}\sin(y)-z\sin(zy)$ and $g^{\prime\prime}(y)=z^{2}\cos(y)-z^{2}\cos(zy)$. Clearly for $y\in[0,\pi/z]$, we have $\cos(zy)\leq\cos(y)$ and thus $g^{\prime\prime}(y)\geq 0$. Note $g^{\prime}(0)=0$ and therefore $g(y)\geq 0$ for $y\in[0,\pi/z]$. Now consider $y\in[\pi/z,\pi]$, $z^{2}(1-\cos(y))\geq z^{2}(1-\cos(\pi/z))\geq 2\geq 1-\cos(zy)$, which implies $g(y)\geq 0$ on $0,\pi]$. Thus we have $g(y)\geq 0$ for $\pi,2\pi]$ too by symmetry. The statement is thus proved since $g(y)$ is periodic in $2\pi$.
For the third statement, $f(z)=0$ necessarily implies that $\sigma=0$ and the measure $\nu$ concentrates on $2\pi j/z$ for $j\in\mathbb{Z}$. Thus $f(x)=\sum_{j\in\mathbb{Z}}(1-\cos(2\pi jx/z))\nu(\{2\pi j/z\})$ is periodic with period $z$. ∎
7.2 II: Tractability over Integers vs. Reals
It is known that $F_{p}$ is intractable for $p>2$ [11, 26, 41, 56]. Nonetheless, there are tractable $f$ that grow slightly super-quadratically, such as $f(x)=x^{2}\log^{c}(1+|x|)$, for $c=O(1)$; see [9, 8]. To be more specific, such $f$ are tractable over the integers, but not the reals, as we now prove.
Theorem 17.
$\mathcal{T}[\mathbb{Z}]\neq\mathcal{T}[\mathbb{R}]$. Specifically, $f(x)=x^{2}\log(1+|x|)\in\mathcal{T}[\mathbb{Z}]-\mathcal{T}[\mathbb{R}]$.
Proof.
When $\mathbf{x}\in\{-n,\ldots,n\}^{n}$ it is known that $(1\pm\epsilon)$-approximating $F_{3}(\mathbf{x})=\sum_{j}|\mathbf{x}(j)|^{3}$ requires $\tilde{\Omega}_{\epsilon}(n^{1/3})$ space [11, 26, 41, 56]. Map $\mathbf{x}$ to $\mathbf{y}=\mathbf{x}/n^{2}$. Then
$f(\mathbf{y})=\sum_{j}f(\mathbf{y}(j))=\sum_{j}\mathbf{y}(j)^{2}\log(1+|\mathbf{y}(j)|)$ $=\sum_{j}\frac{\mathbf{x}(j)^{2}}{n^{4}}\log\left(1+\frac{|\mathbf{x}(j)|}{n^{2}}\right)$
$=\sum_{j}\left(\frac{|\mathbf{x}(j)|^{3}}{n^{6}}-\frac{1}{2}\frac{|\mathbf{x}(j)|^{4}}{n^{8}}+O\left(\frac{|\mathbf{x}(j)|^{5}}{n^{10}}\right)\right)$
$\in[(1-1/n)n^{-6}F_{3}(\mathbf{x}),n^{-6}F_{3}(\mathbf{x})].$
Thus any $(1\pm\epsilon)$-approximation of $f(\mathbf{y})$ yields a $(1\pm(\epsilon+1/n))$ approximation of $F_{3}(\mathbf{x})$. ∎
7.3 III: Nearly Periodic Functions
Braverman, Chestnut, Woodruff, and Yang [8, §5] gave the following example of a nearly periodic function that cannot be solved in their $F_{2}$-heavy hitter-based framework, but can be estimated using ad hoc algorithmic tricks, in $O(\epsilon^{-8}\log^{15}n)$ space.
Definition 12 ($g_{np}$, a nearly periodic function [8, §5]).
For $x\in\mathbb{N}$, $g_{np}(x)=2^{-\tau(x)}$, where $\tau(x)=\max\{j\in\mathbb{N}:2^{j}|x\}$, i.e., the position of the first “1” in the binary representation of $x$.
The reason that $g_{np}$ cannot be tracked by finding $F_{2}$-heavy hitters is that $g_{np}(x)$ can occasionally become polynomially small in $x$. In particular, for $k\in\mathbb{N}$, $g_{np}(2^{k})=2^{-k}$. We now demonstrate how the function $g_{np}$ can be estimated directly using the Lévy-Tower.
To prove that $g_{np}\in\mathcal{T}[\mathbb{Z}]$ we may assume that $\mathbf{x}\in\{-n,\ldots,0,\ldots,n\}^{n}$, so it suffices to consider a $2^{w}$-periodic version of $g_{np}$ with $w>\log n$.
Definition 13 ($2^{w}$-periodic $g_{np}$).
For $x\in\mathbb{N}$,
$g_{np,w}(x)$ $=g_{np}(x\operatorname{mod}2^{w})$
We show that, indeed, $g_{np}$ corresponds to a Lévy process, specifically a compound Poisson process.
Figure 5: Left: $g_{np,5}(x)=\sum_{j=1}^{31}\frac{2^{2\tau(j)+1}+1}{1536}(1-\cos(2\pi jx/32)).$ The black dots mark the values at integer $x$s. Right: The jump distribution of the compound Poisson process $X$ with characteristic exponent $g_{np,5}$. Such nearly periodic functions do not fit in the $F_{2}$-heavy-hitter based framework in [8]. Nevertheless, one may compute the corresponding Lévy process and apply the Lévy-Tower.
Lemma 13.
Fix $w\in\mathbb{N}$. For any $x\in\mathbb{N}$,
$g_{np,w}(x)$ $=\sum_{j=1}^{2^{w}-1}\frac{2^{2\tau(j)+1}+1}{3\cdot 2^{2w-1}}(1-\cos(2\pi jx/2^{w})).$
Proof.
This expression was derived by taking the Fourier transform of $g_{np,w}$ over $\mathbb{Z}/2^{w}\mathbb{Z}$ but we shall prove the lemma without going through Fourier transforms. When $x=0$, $g_{np,w}(x)=0$ as all the cosines are 1. When $x\in(0,2^{w})$ write it as $x=y2^{r}$ where $y$ is odd, so our goal is to show that $g_{np,w}(x)=2^{-r}$.
$g_{np,w}(x)$ $=\sum_{j=1}^{2^{w}-1}\frac{2^{2\tau(j)+1}+1}{3\cdot 2^{2w-1}}(1-\cos(2\pi jx/2^{w}))$
$=\frac{1}{3\cdot 2^{2w-1}}\sum_{s=0}^{w-1}(2^{2s+1}+1)\sum_{\begin{subarray}{c}j=z2^{s}\\ z\text{ odd}\end{subarray}}(1-\cos(2\pi yz/2^{w-r-s}))$
Now when $s\geq w-r$ the second sum vanishes as all cosines are 1. Moreover, $z\in\{1,3,\ldots,2^{w-s}-1\}$ takes on $2^{w-s-1}$ values, and since $y$ is coprime to every power of 2, the second sum runs through the $2^{w-r-s-1}$ odd residues of $2^{w-r-s}$, $2^{r}$ times each. Continuing, we have
$=\frac{1}{3\cdot 2^{2w-1}}\sum_{s=0}^{w-r-1}(2^{2s+1}+1)\cdot 2^{r}\sum_{z\in\{1,3,\ldots,2^{w-r-s}-1\}}(1-\cos(2\pi z/2^{w-r-s}))$
When wrs>1w-r-s>1 the cosines of the odd residues sum to 0, whereas when wrs=1w-r-s=1 there is only one odd residue and $1-\cos(2\pi/2)=2$.
$=\frac{1}{3\cdot 2^{2w-1}}\sum_{s=0}^{w-r-1}(2^{2s+1}+1)\cdot 2^{r}\cdot\left\{\begin{array}[]{ll}2^{w-r-s-1}&\text{ when $w-r-s>1$,}\\ 2&\text{ when $w-r-s=1$.}\end{array}\right$.
$=\frac{1}{3\cdot 2^{2w-1}}\left(\sum_{s=0}^{w-r-2}(2^{2s+1}+1)2^{w-s-1}+(2^{2(w-r-1)+1}+1)2^{r+1}\right)$
$=\frac{1}{3\cdot 2^{2w-1}}\left(2^{w}\sum_{s=0}^{w-r-2}(2^{s}+2^{-s-1})+(2^{2w-r}+2^{r+1})\right)$
$=\frac{1}{3\cdot 2^{2w-1}}\left(2^{w}(2^{w-r-1}-2^{-(w-r-1)})+(2^{2w-r}+2^{r+1})\right)$
$=\frac{3\cdot 2^{2w-r-1}}{3\cdot 2^{2w-1}}=2^{-r}.$
By Lévy-Khintchine, $g_{np,w}$ is the characteristic function of a compound Poisson process $(X_{t})$ with jump rate
$\sum_{j=1}^{2^{w}-1}\frac{2^{2\tau(j)+1}+1}{3\cdot 2^{2w-1}}=\frac{2^{2w}-1}{3\cdot 2^{2w-1}},$
where the jump distribution is symmetric, with jump $J$ distributed as
$\mathbb{P}(J=2\pi j/2^{w})=\mathbb{P}(J=-2\pi j/2^{w})$ $=\frac{1}{2}\frac{2^{2\tau(j)+1}+1}{2^{2w}-1},\quad\text{for $j=1,\ldots 2^{w}-1.$}$
Moreover, the parameter $w$ for the random jump induced by $g_{np,w}$ has a neat algorithmic interpretation as the max recursion depth if one implements each random jump by a recursive algorithm141414(Halt with prob. 1/2; go left and repeat with prob. 1/4; and go right and repeat with prob. 1/4; see Fig. 5).. The case where $w\to\infty$ corresponds to the algorithm without depth limit. Let $\mathbb{B}\subset[0,1)$ be the set of all real fractions with finite binary representations. Note that the jump distribution converges pointwisely to
$\mathbb{P}(J=2\pi x)=\mathbb{P}(J=-2\pi x)$ $=\frac{1}{2}2^{1-2\tau_{*}(x)}=2^{-\tau_{*}(x)},$
for any $x\in\mathbb{B}$, where $\tau_{*}(x)$ is the length of the representation of $x$, e.g., $\tau_{*}(3/8)=\tau_{*}((0.011)_{2})=3$. This jump distribution can be simulated in unbounded input streams as follows. Let $A=\sum_{j=1}^{\infty}A[j]2^{-j}$ be uniform in $[0,1]$, $T\sim\operatorname{Geometric}(1/2)$, and $\xi\in\{-1,1\}$ be Rademacher. Then the jump $J$ is distributed as:
$J\sim 2\pi\,\xi\left(\sum_{j=1}^{T-1}A[j]2^{-j}+2^{-T}\right).$
7.4 IV: The Fourier-Hahn-Lévy Method
The example of $g_{np}$ shows that it is possible to systematically sketch nearly periodic functions by sketching periodic functions and letting the period go to infinity. (See Wang [54] for other examples.) Given the success of the Lévy-Tower in estimating both “standard” $f$-moments and exotic ones like the $g_{np}$-moment, it is natural to conjecture that the Lévy-Khintchine representation theorem actually defines the set of tractable functions $f$. Unfortunately, reality is not quite this clean, and indeed, this simplistic conjecture is absolutely false. We begin by giving a simple function $f$ that is obviously tractable, and obviously not Lévy-Khintchine-representable. Nonetheless, the conjecture can be salvaged by encoding $f$ as the difference between two Lévy-Khintchine-representable functions, via what we call the Fourier-Hahn-Lévy method.
Example 5 (The 0-1-5 Problem).
Consider the function $f:\mathbb{Z}\to\mathbb{R}_{+}$, where
$f(x)=\left\{\begin{array}[]{ll}0&\text{ if $x=0$,}\\ 1&\text{ if $|x|=1$,}\\ 5&\text{ if $|x|>1$.}\end{array}\right$.
The $f$-moment is easy to estimate to within a $1\pm\epsilon$ factor in $O(\epsilon^{-2}\operatorname{polylog}(n))$ space, using $F_{0}$-sampling [19]. However, no Lévy process has its characteristic function equal to $f$, for otherwise we would have $f(2\cdot 1)\leq 2^{2}f(1)$ by Lemma˜12, which implies $5\leq 4$, a contradiction.
7.4.1 The Periodic Fourier-Hahn-Lévy Method
Let us now extend our method to handle the 0-1-5 problem and others. Suppose that $\|\mathbf{x}\|_{\infty}<n$. It suffices to consider a periodic version $f_{p}$ of $f$ with period $p\geq 2n$ that agrees with $f$ on the range $\{-n,\ldots,n\}$. For example, in the case of the 0-1-5 problem we would define $f_{p}$ as
$f_{p}(x)=\left\{\begin{array}[]{ll}0&\text{ if $|x|_{p}=0$,}\\ 1&\text{ if $|x|_{p}=1$,}\\ 5&\text{ if $|x|_{p}>1$,}\end{array}\right$.
where $|x|_{p}=\min\{x\operatorname{mod}p,p-x\operatorname{mod}p\}$. The three-step Fourier-Hahn-Lévy method is executed as follows.
Fourier Transform.
Decompose $f(x)-f(0)=\sum_{j=1}^{p-1}(1-\cos(2\pi jx/p))\hat{f}(j)$ where the coefficient $\hat{f}(j)$ can be computed by taking the Fourier transform of ff. [^15] We have $f(x)-f(0)=\sum_{j=1}^{p-1}\tilde{f}(j)(e^{2\pi ijx/p}-1)$. Since $f$ is real and symmetric $\tilde{f}$ is real and $\tilde{f}(j)=\tilde{f}(p-j)$, so we can pair up the summands $j$ and pjp-j as follows: $\tilde{f}(j)(e^{2\pi ijx/p}-1)+\tilde{f}(p-j)(e^{-2\pi ijx/p}-1)=-2\tilde{f}(j)(1-\cos(2\pi jx/p))=-2\tilde{f}(p-j)(1-\cos(2\pi jx/p))$. Setting $\hat{f}(j)=-\tilde{f}(j)$ we have $f(x)-f(0)=\sum_{j=1}^{p-1}(1-\cos(2\pi jx/p))\hat{f}(j)$. Note that $\hat{f}$ is real since we have assumed $f$ to be real and symmetric.
Hahn Decomposition.
Decompose $\hat{f}=\hat{f}_{+}-\hat{f}_{-}$ where $\hat{f}_{+}$ and $\hat{f}_{-}$ are non-negative real functions.
Lévy Process Simulation.
Define functions $f_{+},f_{-}$ as the inverse Fourier transform of $\hat{f}_{+},\hat{f}_{-}$ respectively. Since the Fourier transform is linear, we have $f=f_{+}-f_{-}$. By construction, $f_{+},f_{-}$ are each characteristic exponents of some Lévy process. Estimate those function moments separately with Lévy-Tower sketches parameterized by $m$, and take their difference to yield an estimate of the $f$-moment.
Lemma 14.
Given any symmetric function $f:\mathbb{Z}\to\mathbb{R}_{+}$ with $f(0)=0$ and any stream $\mathbf{x}\in\mathbb{Z}^{n}$, let $\widehat{f(\mathbf{x})}$ be the estimator derived from the Fourier-Hahn-Lévy method with parameter $p>2\|\mathbf{x}\|_{\infty}$. Then $\widehat{f(\mathbf{x})}$ has error
$\left|\widehat{f(\mathbf{x})}-f(\mathbf{x})\right|=O\left(\frac{|f_{+}(\mathbf{x})|+|f_{-}(\mathbf{x})|}{\sqrt{m}}\right),$
with probability at least 98/100.
Figure 6: Here $f$ is the symmetric, periodic version of the 0-1-5 function over $\mathbb{Z}/p\mathbb{Z}$, namely $f(0)=0,f(1)=f(99)=1,f(x)=5$ for $x\in[2,98]$. It is decomposed into $f=f_{+}-f_{-}$, where in this case $f_{+}+f_{-}=\Theta(f)$.
If $f$ is the characteristic exponent of a Lévy process, then $f=f_{+}$ and $f_{-}=0$. This case was already considered, where we get an asymptotically unbiased estimate with multiplicative error. Clearly, whenever $|f_{+}(\mathbf{x})|+|f_{-}(\mathbf{x})|=O(|f(\mathbf{x})|)$, the Fourier-Hahn-Lévy method still estimates the $f$-moment to within a $(1\pm\epsilon)$-factor, when $m=\Theta(\epsilon^{-2})$. This handles the 0-1-5 problem, among others; see Fig.˜6.
(a) Decomposition of $f=f_{+}-f_{-}$.
(b) Decomposition of $f=f_{+}-f_{-}$ over $[1,3]$.
(c) Ratio $(f_{+}+f_{-})f$ over $[1,3]$.
Figure 7: The (periodic) Fourier-Hahn-Lévy method applied to $f(x)=x^{2}\log(1+|x|)$ over the domain [100,100][-100,100]. Ther periodic version of $f$ is over $\mathbb{Z}/200\mathbb{Z}$.
Let us explore how the Fourier-Hahn-Lévy method performs on functions that grow slightly faster than quadratic, say $f(x)=x^{2}\log(1+|x|)$. Fig.˜7 illustrates the decomposition of the periodic form $f_{p}(x)=|x|_{p}^{2}\log(1+|x|_{p})$, where $n=100,p=200$, which agrees with $f$ on $\{-n,\ldots,n\}$. This decomposition has an impractically large ratio $\max_{x\in[1,n]}(f_{+}(x)+f_{-}(x))/f(x)$, which is over 1,4001{,}400 when $x=1$.
A better approach uses the fact that real Lévy-Khintchine-representable functions are of the form $Ax^{2}+\int_{0}^{\infty}(1-\cos(xs))\nu(ds)$. When $|x|\leq n,p=2n$, we may write $f(x)=x^{2}\log(1+n)-g_{p}(x)$, where $g_{p}(x)=|x|_{p}^{2}\log(\frac{1+n}{1+|x|_{p}})$ is non-negative over the relevant range $|x|\leq n$, and $x^{2}\log(1+n)$ is already Lévy-Khintchine-representable, $\log(1+n)$ being constant. Applying the periodic Fourier-Hahn-Lévy method to $g=g_{p}$ over $\mathbb{Z}/p\mathbb{Z}$, yields $g=g_{+}-g_{-}$, where $g_{+},g_{-}$ are Lévy-Khintchine-representable. Since Lévy-Khintchine-representable functions are closed under addition we can decompose $f=f_{+}-f_{-}$ into Lévy-Khintchine-representable functions such that whenever $|x|\leq n$,
$f(x)=x^{2}\log(1+n)-g(x)=(x^{2}\log(1+n)+g_{-}(x))-g_{+}(x)=f_{+}(x)-f_{-}(x),$
where $f_{+}(x)=x^{2}\log(1+n)+g_{-}(x)$ and $f_{-}(x)=g_{+}(x)$. Note that $g_{-}$ and $g_{+}=f_{-}$ are periodic, whereas $f_{+}$ is not. As one can see in Fig.˜8 this mixed-periodic decomposition has a substantially better ratio $(f_{+}(x)+f_{-}(x))/f(x)$, but it is still impractical.
(a) Decomposition of $g=g_{+}-g_{-}$.
(b) Decomposition of $f=f_{+}-f_{-}$.
(c) Decomposition of $f=f_{+}-f_{-}$ over $[1,3]$.
(d) Ratio $(f_{+}+f_{-})f$.
Figure 8: An improved Fourier-Hahn decomposition of $f(x)=x^{2}\log(1+|x|)$ over $|x|\leq n$, $n=100$. We write $f(x)=x^{2}\log(1+n)-g(x)$, and apply the Fourier-Hahn-Lévy method to decompose $g$ over $\mathbb{Z}/(2n)\mathbb{Z}$.
7.4.2 The Approximate Fourier-Hahn-Lévy Method
At this point we realize that it is counterproductive to consider only exact decompositions of $f=f_{+}-f_{-}$. In order to $(1\pm\epsilon)$-approximate the $f$-moment $f(\mathbf{x})$ with a $\operatorname{poly}(\epsilon^{-1},\log n)$-bit sketch, it suffices to find a function $g=g_{+}-g_{-}$ such that $g_{+},g_{-}$ are Lévy-Khintchine-representable and for all relevant $|x|<n$, $g(x)\in[(1-\epsilon)f(x),(1+\epsilon)f(x)]$ is a good approximation of $f$ and $g_{+}(x)+g_{-}(x)\leq\operatorname{poly}(\epsilon^{-1},\log n)f(x)$ is not too much larger than $f$. (It is not necessary that $g$ be periodic, nor that it have any prescribed behavior outside of $|x|<n$.) Rather than do this analytically, we can generate $g,g_{+},g_{-}$ by solving a linear program encoding the relevant constraints. Let $(\omega_{k})$ be a sufficiently dense set of frequencies. For approximating $f$, $\omega_{k}=(1+\epsilon)^{-k}$ for $k\in\{-\Omega(\epsilon^{-1}),\ldots,0,\ldots,\log_{1+\epsilon}n+O(1)\}$ suffices. Let $u_{k},v_{k}$ be the non-negative coefficients of $g_{+},g_{-}$ corresponding to $\omega_{k}$, and $u_{*},v_{*}$ be the coefficients of $x^{2}$. In other words, $g_{+},g_{-}$ are
$g_{+}(x)=u_{*}x^{2}+\sum_{k}u_{k}(1-\cos(\omega_{k}x))\qquad\text{and}\qquad g_{-}(x)=v_{*}x^{2}+\sum_{k}v_{k}(1-\cos(\omega_{k}x)).$
The linear program is then:
Minimize $\rho$ subject to:
$\forall x\in\{1,\ldots,n\}.\quad(u_{*}-v_{*})x^{2}+\sum_{k}(u_{k}-v_{k})(1-\cos(\omega_{k}x))$ $\in[(1-\epsilon)f(x),(1+\epsilon)f(x)],$
$\forall x\in\{1,\ldots,n\}.\quad(u_{*}+v_{*})x^{2}+\sum_{k}(u_{k}+v_{k})(1-\cos(\omega_{k}x))$ $\leq\rho f(x),$
$u_{k},v_{k}$ $\geq 0.$
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Refer to caption Refer to caption
(e) (f)
Refer to caption Refer to caption
(g) (h)
Figure 9: Plots (a)–(d) present the results of the LP construction of $g_{+},g_{-}$ with $n=1000,\epsilon=0.03$ while plots (e)–(h) are for $n=1000,\epsilon=0.003$. Plots (d,h) show that the cost of the decomposition $(g_{+}(x)+g_{-}(x))/f(x)\leq\rho$ does not depend heavily on $\epsilon$, for $\epsilon$ in this range.
Figure 9 illustrates the results of this LP when $n=1000,\epsilon=0.03$, plots (a)–(d), and $n=1000,\epsilon=0.003$, plots (e)–(h). As one can see, the ratio $\rho=\max_{x\in[1,n]}(g_{+}(x)+g_{-}(x))/f(x)$ of these solutions (plots (d,h)) are much more reasonable than the (exact) applications of the periodic or mixed-periodic Fourier-Hahn-Lévy method.
8 Conclusion
In this paper we introduced a new way to study the tractability question in data sketches, which has historically been studied alongside the universality question [9, 8, 7, 54].
Tractability Question.
What is the class $\mathcal{T}$ of $f$-moments that can be $(1\pm\epsilon)$-estimated in $\operatorname{poly}(\epsilon^{-1},\log n)$ space?
Universality Question.
A sketch is $\mathcal{C}$-universal if it can $(1\pm\epsilon)$-estimate the $f$-moment, for any $f\in\mathcal{C}$. Is there a $\mathcal{T}$-universal sketch?
We have demonstrated that the Lévy-Tower parameterized by a suitable Lévy process $(X_{t})$ can estimate the $f_{X}$-moment, where $f_{X}$ is the characteristic exponent of $X$. Moreover, many existing sketches can be reinterpreted as sampling from Lévy processes. At first glance this approach seems suited to exploring the tractability question, but ill-suited to the universality question, for how can the Lévy-Tower(X)(X) be $\mathcal{T}$-universal if it is only built for estimating a single $f_{X}$-moment? An unexpected outcome of this work is the revelation that some Lévy processes $(X_{t})$ “leak” lots of information about $f$-moments unrelated to $f_{X}$. In particular, when a characteristic exponent $f:\mathbb{R}\to\mathbb{R}^{+}$ is real, it can be expressed as $f(x)=Ax^{2}+\int_{0}^{\infty}(1-\cos(xs)\nu(ds))$ for some real $A$ and measure $\nu$; see Remark˜1, and Sections˜6.5 and 7.3 for two examples. In a companion paper by the second author, Wang [54] demonstrates that a data structure similar to Lévy-Tower parameterized by a symmetric Poisson process (SymmetricPoissonTower) can estimate $F_{2}$-moments, all the basis moments $(f_{s})_{s>0}$, where $f_{s}(x)=1-\cos(sx)$, and any linear mixture of the basis moments. This sketch is universal for all the usual tractable real moments, and “natively” handles nearly periodic $f$-moments not amenable to $F_{0}$-sampling or $F_{2}$-heavy hitter sampling [9, 8, 7].
8.1 Conjectures
Is the Lévy-Khintchine representation theorem the final answer to the fundamental tractability [9, 8, 7] question? In Section˜7.2 we muddied the waters a bit by showing that $\mathcal{T}[\mathbb{Z}]$ and $\mathcal{T}[\mathbb{R}]$ are different classes, i.e., tractability with respect to $\mathbf{x}\in\mathbb{Z}^{n}$ vs. $\mathbf{x}\in\mathbb{R}^{n}$ are two different concepts. In Section˜7.4 we showed there exist simple, tractable functions with no Lévy-Khintchine-representation, then demonstrated that such functions can still be handled in the Lévy-Tower/Lévy-Khintchine framework via the periodic Fourier-Hahn-Lévy method or the more general approximate Fourier-Hahn-Lévy method. We believe the Fourier-Hahn-Lévy methods are sufficiently powerful to capture all tractability classes $\mathcal{T}[M]$, where $\mathbf{x}\in M^{n}$ and $M$ is any of the domains of interest, $M\in\{\mathbb{Z},\mathbb{R},\mathbb{Z}^{d},\mathbb{R}^{d}\}$. We state this conjecture formally for $M=\mathbb{Z}$ or $\mathbb{R}$.
Conjecture 1 (Tractability).
Suppose that $f\in\mathcal{T}[\mathbb{Z}]$ (respectively, $f\in\mathcal{T}[\mathbb{R}]$). For any $\epsilon>0$ and parameter $n$ controlling the length of $\mathbf{x}\in\mathbb{Z}^{n}$ (respectively, $\mathbb{R}^{n}$) and $\|\mathbf{x}\|_{\infty}<n$, there exists a $g=g_{+}-g_{-}$ satisfying the following criteria:
  • $g_{+}$ and $g_{-}$ are Lévy-Khintchine-representable,
  • $g(x)\in[(1-\epsilon)f(x),(1+\epsilon)f(x)]$ for all $x\in\{-n,\ldots,0,\ldots,n\}$ (respectively, $x\in[-n,n]$), and
  • $(g_{+}(x)+g_{-}(x))/f(x)=O(\operatorname{poly}(\epsilon^{-1},\log n))$.
It is possible that the connection between space-efficient $G$-samplers and the Laplace exponents of subordinators161616(non-negative, one-dimensional Lévy processes) is also natural, but here we have to be much more stringent in our space requirements to make a plausible conjecture.
Conjecture 2 (Sampling).
Suppose there is a $O(\log n)$-bit data structure for $G$-sampling from a vector $\mathbf{x}\in\mathbb{N}^{n}$ subject to incremental updates. Then $G$ is the Laplace exponent of a non-negative Lévy process on $\mathbb{R}_{+}\cup\{\infty\}$.
The key criterion of Conjecture˜2 is the $O(\log n)$ space bound, which permits techniques like storing the minimum hash value and index, but not more sophisticated techniques [42, 2, 36]. The conjecture would be false if we loosened the space bound to $\operatorname{polylog}(n)$ bits, as $F_{p}$ sampling is possible within this bound for $p\in[0,2]$ [31] but not in correspondence with a subordinator when $p\in(1,2]$ [50].
Appendix A Sampling Without Replacement
We can take $k$ independent copies of the Lévy-Min-Sampler or ParetoSampler sketches to sample $k$ indices from the $(G(\mathbf{x}(v))/G(\mathbf{x}))_{v\in[n]}$ distribution with replacement. A small change to these algorithms will sample $k$ indices without replacement. See Cohen, Pagh, and Woodruff [17] for an extensive discussion of why WOR (without replacement) samplers are often more desirable in practice. The algorithm $(G,k)$-Sampler-WOR(Algorithm˜6) samples $k$ (distinct) indices without replacement.
Sketch : $S\subset[n]\times\mathbb{R}_{+}$, initialized as $\emptyset$
Hash function : $H:[n]\to\mathrm{Uniform}(0,1)$
Result: Sample $k$ elements $u$ with prob. $G(\mathbf{x}(u))/\sum_{v\in[n]}G(\mathbf{x}(v))$ without replacement
// upon update $\mathbf{x}(v)\leftarrow\mathbf{x}(v)+\Delta$
Update $(v,\Delta)$
 $S\leftarrow k\textsf{-Min}(S\cup\{(v,\ell_{G}(\mathrm{Exp}(\Delta),H(v)))\})$
 // $\mathrm{Exp}(\Delta)$ is freshly sampled
// upon sample
Sample ()(\,)
 return $\{v:(v,\cdot)\in S\}$
Algorithm 6 Lévy-Min-Sampler without replacement. The function $k\textsf{-Min}(L)$ takes a list $L\subset[n]\times\mathbb{R}_{+}$, discards any $(v,h)\in L$ if there is a $(v,h^{\prime})\in L$ with $h^{\prime}<h$, then returns the $k$ elements with the smallest second coordinate.
In a similar fashion, one can define a ParetoSampler without replacement that maintains the minimum $k$-Pareto frontier, defined by discarding any tuple $(a,b,v)$ if there is another $(a^{\prime},b,v)$ with $a^{\prime}<a$, then retaining only those tuples that are dominated by at most k1k-1 other tuples.
Theorem 18.
Consider a stream of $\operatorname{poly}(n)$ incremental updates to a vector $\mathbf{x}\in\mathbb{R}_{+}^{n}$. The $(G,k)$-Sampler-WOR occupies $2k$ words of memory, and can report an ordered tuple $(v_{*}^{1},\ldots,v_{*}^{k})\in[n]^{k}$ such that
$\mathbb{P}((v_{*}^{1},\ldots,v_{*}^{k})=(v^{1},\ldots,v^{k}))=\prod_{i=1}^{k}\frac{G(\mathbf{x}(v^{i}))}{G(\mathbf{x})-\sum_{j=1}^{i-1}G(\mathbf{x}(v^{j}))}.$ (11)
The $k$-ParetoSampler occupies $O(k\log n)$ words w.h.p. and for any $G\in\mathcal{G}$ at query time, can report a tuple $(v_{*}^{1},\ldots,v_{*}^{k})\in[n]^{k}$ distributed according to Eq.˜11.
Proof.
The proof of Theorem 13 shows that $h_{v}\sim\mathrm{Exp}(G(\mathbf{x}(v)))$ and if $v_{*}^{1}$ minimizes $h_{v_{*}^{1}}$, that $h_{v_{*}^{1}}\sim\mathrm{Exp}(G(\mathbf{x}))$. It follows that $\mathbb{P}(v_{*}^{1}=v)=G(\mathbf{x}(v))/G(\mathbf{x})$. By the memoryless property of the exponential distribution, for any $v\neq v_{*}^{1},$ $h_{v}-h_{v_{*}^{1}}\sim\mathrm{Exp}(G(\mathbf{x}(v)))$, hence $h_{v_{*}^{2}}-h_{v_{*}^{1}}\sim\mathrm{Exp}(G(\mathbf{x})-G(\mathbf{x}(v_{*}^{1})))$ and $\mathbb{P}(v_{*}^{2}=v\mid v_{*}^{1},v\neq v_{*}^{1})=G(\mathbf{x}(v))/(G(\mathbf{x})-G(\mathbf{x}(v_{*}^{1})))$. The distribution of $v_{*}^{3},\ldots,v_{*}^{k}$ is analyzed in the same way.
By the 2D-monotonicity property, the $k$-Pareto frontier contains all the points that would be returned by $(G,k)$-Sampler-WOR, hence the output distribution of $k$-ParetoSampler-WOR is identical. The analysis of the space bound follows the same lines, except that $X_{i}$ is the indicator for the event that $h_{v_{i}}$ is among the $k$-smallest elements of $\{h_{v_{1}},\ldots,h_{v_{i}}\}$, so $\mathbb{E}(X_{i})=\min\{k/i,1\}$, $\mathbb{E}(|S|)<kH_{n}$, and by a Chernoff bound, $|S|=O(k\log n)$ with high probability. ∎